### 30 Mar 2026 v1

# Probing soft signals of gravitational-wave memory with space-based interferometers

Abstract
Gravitational-wave displacement memory is a remarkable and ubiquitous phenomenon predicted by general relativity, which has not yet been detected. Unlike the oscillatory components of gravitational waveforms, displacement memory is associated with soft gravitons, making it the only observable signal of its parent event at sufficiently low frequencies. Similarly, soft waveforms may arise from velocity and integrated-displacement memory. The simple and universal spectral shapes of soft waveforms also provide effective templates for matched filtering and parameter estimation. In this paper, we systematically investigate the detection prospects for such soft memory signals with future space-based laser interferometers. As realistic examples, we examine the infrared spectral features of gravitational waves from moderately relativistic compact binary scattering and quasi-circular, non-precessing black hole mergers. In both cases, the low-frequency spectrum can be described by a soft waveform of displacement memory with a real correction factor. The results of simulated Bayesian parameter estimation demonstrate that independent measurement of a soft displacement-memory signal with a single LISA-like detector is achievable at signal-to-noise ratios $\gtrsim 10$. The measurement precision can be significantly improved by joint observations with a LISA-Taiji network. A single BBO detector would be capable of separately measuring the null memory from stellar-mass compact binary mergers. We also evaluate the detectability of an idealized stochastic background of soft displacement-memory signals. Our results indicate that gravitational-wave bursts with memory can be promising targets for space-based interferometers.
I Introduction
Besides the oscillatory components, gravitational radiation in general relativity generally contains a non-oscillatory component that results in a displacement memory [10, 18, 85, 31]. It may have an “ordinary” or “null” origin [8], and is associated with the supertranslation charge balance laws of asymptotically flat spacetimes [91, 33, 78, 67] as well as with Weinberg’s soft graviton theorem [78]. Displacement memory can therefore serve as an important test of general relativity. In addition to their theoretical interest, gravitational wave (GW) bursts with displacement memory are expected to be produced by a variety of astrophysical sources, including compact binaries [30, 29, 32, 84, 94, 26, 87, 63, 100, 55, 47, 96, 24, 44, 64, 38, 48, 28, 22, 5], gamma-ray bursts [75, Yu_2020, 76], and core-collapse supernovae [68, 72]. Searches for displacement memory in the GWs from compact binary mergers have been considered for ground-based detectors [62, 66, 94, 27, 53, 52, 87, 42, 17, 41], space-based detectors [56, 40, 39, 54, 81, 41, 19, 101], and pulsar timing arrays [88, 89, 56, 2, 79, 25, 80, 1]. Potentially observable imprints of displacement memory may also appear in the cosmic microwave background [102].
Although displacement memory ubiquitously occurs in gravitational radiation, it has so far remained undetected. For example, the null memory of a compact binary merger produces only a small modification of the waveform and is not easily disentangled from the oscillatory “parent” signal in the time domain. Resolving such modifications through sensitive detectors and refined waveform modeling would enable the detection of memory and the improvement of parameter estimation [42, 39, 54, 81]. Accurately modeling waveforms that include memory effects becomes increasingly challenging for more complex sources. For a compact binary scattering event, even with a sufficiently accurate waveform model, parameter estimation is still challenging due to degeneracies [fontbuté2024numericalrelativitysurrogatemodelhyperbolic]. Waveform modeling is further complicated by possible environmental influences [15, 77]. However, because of its soft nature, displacement memory produces a signal that persists at low frequencies, with a spectrum of universal shape. Although much information about the sources is lost in such “orphan” signals [88, 66, 40], they remain valuable targets for probing the memory components—either by searching for them or by measuring them independently—thanks to the ubiquity of the sources and the simplicity of the templates. It is also possible that some yet-unknown processes could produce observable GW bursts with displacement memory, making soft memory signals a potential probe of these processes.
In addition to the soft waveform associated with displacement memory (which corresponds to a step-like strain in the time domain), other types of soft waveforms may also be observed, including those associated with velocity memory [43, 9] (corresponding to a step-like first derivative of strain in the time domain) and integrated-displacement memory (discussed in Sec. II.3). Although these waveforms are not related to asymptotic symmetries or the soft theorem, they can arise over sufficiently long time scales and are likewise associated with soft gravitons.
In this paper, we study the prospects for detecting and measuring such soft memory signals with future space-based interferometers, including Laser Interferometer Space Antenna (LISA) [3, 4], Taiji [50], TianQin [50], and Big Bang Observer (BBO) Stage I [20]. We investigate the detector responses and the parameter estimation precision for both a single LISA-like detector and the proposed LISA-Taiji network [74]. In optimal configurations, the signal-to-noise ratio in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$, or a velocity memory of amplitude $\sim 10^{-22}\,\text{Hz}$. To examine the precision of the measurement, we perform numerical simulations for Bayesian parameter estimation. We find that with a single detector, the memory parameters can be effectively probed for signal-to-noise ratios $\gtrsim 10$, although the measurement precision is limited by the signal’s degeneracy at low frequencies. This limitation can be mitigated by using a detector network or by incorporating prior information on the source direction. We demonstrate that the joint observations by LISA and Taiji can achieve significantly more precise measurements. Soft memory signals in BBO (occurring at $f\lesssim 1\,\text{Hz}$) appear to suffer from less degeneracy and can achieve much higher sensitivity, with the signal-to-noise ratio reaching 10 for a displacement memory of amplitude $\sim 10^{-24}$, or a velocity memory of amplitude $\sim 10^{-23}\,\text{Hz}$. This makes it possible to independently measure the null displacement memory from stellar-mass compact binary mergers.
The soft waveform is an idealized zero-frequency limit. We examine the correction to the soft waveform in two realistic examples: hyperbolic encounter of compact binary and quasi-circular BBH merger, and assess its impact on displacement memory measurements using the uncorrected template. We find that, in the considered examples, the memory amplitude can still be effectively constrained compared to the case without correction. Soft signals may also contribute to the stochastic gravitational-wave background. We simulate an idealized background of soft displacement-memory signals, and evaluate its detectability at the considered detectors. Our results highlight soft signals from bursts with memory as potential targets for space-based gravitational-wave detectors.
This paper is organized as follows. In Sec. II, we introduce the soft waveforms and discuss their main properties. Sec. III examines the low-frequency GW spectrum of compact binary scattering and merger events, as concrete examples of realistic waveforms exhibiting displacement memory. In Sec. IV, we investigate the responses of space-based detectors to soft memory signals. Sec. V presents the results of simulated parameter estimation. The stochastic background of soft displacement-memory signals and its detectability in space-based detectors is discussed in Sec. VI. We provide a brief summary and discussion in Sec. VII. Throughout this paper, we use the natural units $c=G=\hbar=1$ and the mostly-plus metric signature.
II Soft memory waveforms
We refer a specific form of the zero-frequency limit of the spectrum that depends only on the asymptotic difference in a waveform observable in the time domain as a soft waveform. This is a useful notion, since such a limit is shared by all waveforms with the same “memory” and becomes a good approximation to the actual signal at sufficiently low frequencies. In this section, we introduce three types of soft waveforms, corresponding respectively to the displacement memory, velocity memory and integrated-displacement memory. In the subsequent sections on detectability, however, we focus primarily on displacement memory.
The GWs emitted by distant sources are well-approximated by plane waves when they reach our detectors. Consider a flat spacetime perturbed by a weak plane GW in the transverse-traceless (TT) gauge:
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-dt^{2}+(\delta_{ij}+h_{ij})\,dx^{i}dx^{j},$ (1)
where $h_{ij}(t,\mathbf{x})$ is the strain tensor field satisfying $\delta_{ij}h_{ij}=\partial_{i}h_{ij}=0$. The strain tensor can be decomposed into two linearly-polarized states, and the waveform at $\mathbf{x=0}$ reads
$h_{ij}(t)\equiv h_{ij}(t,\mathbf{0})=\sum_{\lambda=+,\times}h_{\lambda}(t)\,e_{ij}^{\lambda}.$ (2)
Here the ortho-normalized polarization tensors are constructed as
$e_{ij}^{+}=a_{i}a_{j}-b_{i}b_{j},\quad e_{ij}^{\times}=a_{i}b_{j}+b_{i}a_{j}.$ (3)
with the three unit vectors $\{\mathbf{a,b,\hat{k}}\}$ forming a right-handed triad, such that the GW propagating along the direction of $\hat{\mathbf{k}}$, and $e_{ij}^{\lambda}e_{ij}^{\lambda^{\prime}*}=2\delta_{\lambda,\lambda^{\prime}}$. Now we consider the waveform of a given linearly-polarized state, temporarily denoted as $h(t)$ in this section. For the Fourier transform, we adopt the definition:
$h(t)$ $=\int_{-\infty}^{\infty}df\,e^{-2\pi ift}\,\tilde{h}_{\infty}(f),$ (4)
$\tilde{h}_{\infty}(f)$ $\equiv\int_{-\infty}^{\infty}dt\,e^{2\pi ift}\,h(t),$ (5)
where the subscript $\infty$ is a reminder of the idealistic assumption of an infinite time duration. Since $h(t)\in\mathbb{R}$, $\tilde{h}_{\infty}(-f)=\tilde{h}_{\infty}^{*}(f)$.
To see the intuitive observable effects of displacement and velocity memory, we introduce a free-falling timelike test particle at $\mathbf{x=0}$ in the spacetime (1) and the Fermi normal coordinates (FNC) $\bar{x}^{a}=(\bar{t},\bar{\mathbf{x}})$ centered at it, with $\frac{\partial x^{\mu}(\bar{x})}{\partial\bar{x}^{a}}\approx\delta^{\mu}_{a}$ and the particle’s proper time $\bar{t}\approx t$. In this FNC, the equation of motion for a nearby free-falling classical test particle with $\bar{x}^{j}(\bar{t}=0)=X^{j}$ and negligible initial velocity is
$\frac{d^{2}\bar{x}^{i}}{d\bar{t}^{2}}\approx\frac{1}{2}\,\ddot{h}_{ij}(\bar{t})\,\bar{x}^{j}\approx\frac{1}{2}\,\ddot{h}_{ij}(\bar{t})\,X^{j}.$ (6)
The response of an interferometer to low-frequency GWs can be understood as arising from such relative displacements between test masses in the FNC, where the effects of GWs on light propagation are negligible. However, this approximation is not adequate for space-based detectors, and (as discussed in Sec. III) the signals must be analyzed in TT coordinates.
II.1 Displacement memory
Displacement memory refers to a DC (direct current) shift of $h(t)$ between the asymptotic past and future, i.e., $\Delta h\equiv h(\infty)-h(-\infty)\neq 0$. In Eq. (6), this leads to a “permanent” displacement: $\Delta\bar{x}^{i}\approx\frac{1}{2}\Delta h_{ij}X^{j}$. In practice, $\infty$ refers to a moment that is so much far away from the observer’s time of interest. The soft waveform of displacement memory depends solely on the height $\Delta h$ of the memory jump. As illustrative examples, we consider three simple models with exactly computable frequency spectra. The first one is the arctan model:
$h(t)$ $=\frac{\Delta h}{\pi}\arctan\left(f_{*}(t-t_{*})\right),$ (7)
$\tilde{h}_{\infty}(f)$ $=\frac{i\Delta h}{2\pi f}e^{-2\pi|f|/f_{*}}e^{2\pi ift_{*}},$ (8)
with $f_{*}>0$. Note that here for the Fourier transform we are treating $h(t)$ as a tempered distribution, and a global DC shift of the time-domain waveform is unobservable in the spectrum at $f>0$. The jump arises exactly from the zero-frequency mode, which can be seen by imposing an infrared cutoff at $f_{c}>0$. For the waveform (8), one has
$\int_{f_{c}}^{\infty}df\,e^{-2\pi ift}\tilde{h}_{\infty}(f)=\frac{i\Delta h}{2\pi}\Gamma\left(0,2\pi f_{c}(if_{*}t+1)/f_{*}\right),$ (9)
where $\Gamma(a,z)=\int_{z}^{\infty}dt\,t^{a-1}e^{-t}$ is the incomplete Gamma function. This contribution vanishes at $t\to\pm\infty$. A similar example is the tanh model: $h(t)=\frac{\Delta h}{2}\tanh\left(f_{*}(t-t_{*})\right)$, with $\tilde{h}_{\infty}(f)=\frac{i\pi\Delta h/f_{*}}{2\sinh\left(f\pi^{2}/f_{*}\right)}e^{2\pi ift_{*}}$. A linear growth in $t\in t_{*}+[0,1/f_{*}]$ provides another analytically tractable waveform with displacement memory: $h(t)/\Delta h=f_{*}(t-t_{*})\,\Theta(t-t_{*})\,\Theta(t_{*}+1/f_{*}-t)+\Theta(t-t_{*}-1/f_{*})$, with $\tilde{h}_{\infty}(f)=\Delta h\left[\frac{e^{2\pi if/f_{*}}-1}{4\pi^{2}f^{2}/f_{*}}+\frac{\delta(f)}{2}\right]e^{2\pi ift_{*}}$, where $\Theta(t)$ is the Heaviside unit step function.
In all three models above, $1/f_{*}$ measures the time scale of the jump. In the limit $f_{*}\to\infty$, the waveform tends to a step at $t=t_{*}$, with the height $\Delta h$. For a finite $f_{*}$, the models have the same zero-frequency limit at $f/f_{*}\to 0^{+}$:
$\tilde{h}_{\infty}(f)\to\frac{i\Delta h}{2\pi f}\,e^{2\pi ift_{*}},$ (10)
since the details of the jump are unresolvable in this limit. Note that the spectral density of the graviton number flux associated with this waveform diverges at zero frequency, but that of the energy flux remains finite. (10) is shared by all waveforms with the same displacement memory $\Delta h$. For $f\ll 1/t_{*}$, the phase factor $e^{2\pi ift_{*}}$ can be further approximated by unity. This however is generally a bad approximation for the GW signal, since the frequency spectrum of the waveform can only be measured at finite frequency, with the lowest resolvable frequency limited by the signal duration and instrumental noise. By the same reason, detection of the soft waveform (10) cannot directly prove the existence of the memory jump (which is associated with zero-frequency gravitons), but will provide a decisive evidence. Likewise, a detector capable of observing the spectrum in the form (10) at frequency $f>f_{\text{min}}$ has no information about it at lower frequencies. A more complete measurement of the waveform can thus only rely on the detectors having better sensitivities at low frequencies.
We note that for a waveform $h(t)$ with displacement memory satisfying $h(t<t_{\text{min}})\approx h(-\infty)$ and $h(t>t_{\text{max}})\approx h(\infty)$, $\tilde{h}_{\infty}(f)$ is drastically different from the Fourier transform of $h(t)$ in the finite interval $t_{\text{min}},t_{\text{max}}]$:
$\tilde{h}(f)$ $\equiv\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}\,h(t)$
$=\tilde{h}_{\infty}-\frac{i}{2\pi f}\left[h(\infty)\,e^{2\pi ift_{\text{max}}}-h(-\infty)\,e^{2\pi ift_{\text{min}}}\right],$
which in particular is dominated by the second term at high frequencies. For a GW interferometer, however, the observable is not $h(t)$, but its temporal difference: h(tD)h(t)h(t-D)-h(t), with $D$ being constant within a sufficiently short time. So long as $h(t_{\text{min}}+\text{min}(0,-D))\approx h(-\infty)$ and $h(t_{\text{max}}+\text{max}(0,-D)\approx h(\infty)$, because
$\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}[h(t-D)-h(t)]$ (11)
$\qquad\approx\int_{-\infty}^{\infty}dt\,e^{2\pi ift}[h(t-D)-h(t)]$
$\qquad=\left(e^{2\pi ifD}-1\right)\tilde{h}_{\infty}(f),$
the frequency spectrum of the signal well below the sampling frequency can be approximately derived with $\tilde{h}_{\infty}(f)$, but not $\tilde{h}(f)$. Numerically, $\tilde{h}_{\infty}(f)$ can be simply obtained from $\dot{h}(t)$ as $\tilde{\dot{h}}_{\infty}/(-2\pi if)$, with $\tilde{\dot{h}}_{\infty}\approx\tilde{\dot{h}}$ well-approximated by its discrete Fourier transform (DFT). Another way of computing $\tilde{h}_{\infty}(f)$ is to decompose $h(t)$ as a simple waveform with memory such as the arctan model, and a remaining part that has no memory; the latter is then well-approximated by its DFT.
If the time interval $t_{\text{min}},t_{\text{max}}]$ is too short to include the entire jump, the observed signal will be quite different from the soft waveform. As an extreme example, consider a linearly growing waveform $\dot{h}(t)=\text{const}$, for which $h(t-D)-h(t)\propto D$ is approximately constant, and will actually be removed (as a DC component) in the data processing. Moreover, in this paper, we focus on the measurements using space-based interferometers. As explained in Sec. III, the relevant time-domain signals (e.g., Eq. (38)) are built up by observables such as $h(t-D_{1})-h(t)]-[h(t-D_{1}-D_{2})-h(t-D_{2})]$, with $D_{1,2}$ being constant within a sufficiently short time. A linearly growing waveform thus has a vanishing signal in the static-detector approximation. In fact, we find that the signal remains extremely weak compared with that of a soft memory waveform with $\Delta h=\dot{h}\,T$ when the signal duration $T$ is not short, so that the static-detector approximation no longer holds.
II.2 Velocity memory
Analogues to the displacement memory, the velocity memory is a DC shift of $\dot{h}(t)$: $\Delta\dot{h}\equiv\dot{h}(\infty)-\dot{h}(-\infty)\neq 0$. In Eq. (6), this leads to a “permanent” velocity change: $\Delta\frac{d\bar{x}^{i}}{d\bar{t}}\approx\frac{1}{2}\Delta\dot{h}_{ij}X^{j}$. Since $\lim_{f\to 0}\tilde{\dot{h}}_{\infty}(f)\to\frac{i\Delta\dot{h}}{2\pi f}e^{2\pi ift_{*}}$, the soft waveform is given by
$\tilde{h}_{\infty}(f)\to-\frac{\tilde{\dot{h}}_{\infty}}{2\pi if}=-\frac{\Delta\dot{h}}{(2\pi f)^{2}}e^{2\pi ift_{*}}.$ (12)
As mentioned above, we are interested in observables such as $h(t-D_{1})-h(t)]-[h(t-D_{1}-D_{2})-h(t-D_{2})]$, which tends to zero at $t\pm\infty$. The frequency spectrum of the signal can thus also be approximately obtained with $\tilde{h}_{\infty}(f)$, if the observation time is sufficiently long.
A simple example of the waveform with velocity memory is the ReLU (Rectified Linear Unit) model: $h(t)=\Delta\dot{h}_{+}\,\text{ReLU}(t-t_{*})+\Delta\dot{h}_{-}\,\text{ReLU}(t_{*}-t)$, with $\text{ReLU}(t)=t\,\Theta(t)$ being the time integration of $\Theta(t)$. The frequency domain waveform at $f>0$ is exactly given by Eq. (12), with $\Delta\dot{h}=\Delta\dot{h}_{+}+\Delta\dot{h}_{-}$.
II.3 Integrated-displacement memory
It is tempting to consider a waveform with $\frac{d^{n}h}{dt^{n}}(\infty)-\frac{d^{n}h}{dt^{n}}(-\infty)\neq 0$, with $n\geq 2$. However, such a waveform corresponds to a Riemann tensor $R_{0i0j}\approx-\frac{1}{2}\partial_{t}^{2}h_{ij}$ with a constant or polynomially growing amplitude at $t\to\pm\infty$ and is less likely to be associated with a GW [43]. Soft waveforms of the form $\frac{O}{|f|}$ and $\frac{iO}{f|f|}$ with $O\in\mathbb{R}$ are unphysical, since they correspond to time-domain waveforms that grow as $\ln|t|$ and $t\ln|t|$, respectively, as $t\to\pm\infty$.
If hereditary observables are taken into account, additional types of soft waveforms can be found. For example, the waveform $h(t)$ with a finite zeroth time moment is accompanied with an asymptotic change in the observable $O(t)\equiv\int_{-\infty}^{t}dt^{\prime}\,h(t^{\prime})$, and has the zero-frequency limit:
$\tilde{h}_{\infty}(f)\to O\,e^{2\pi ift_{*}},$ (13)
with $O=O(\infty)$. We thus call this the soft waveform of an “integrated-displacement memory”. Note that the imaginary version of (13): $\lim_{f\to 0}\tilde{h}_{\infty}(f)\to iO\,\text{sgn}(f)\,e^{2\pi ift_{*}}$ with $O\in\mathbb{R}$, corresponds to an asymptotic time-domain waveform $\lim_{|t|\to\infty}h(t)\to\frac{O}{\pi}\frac{1}{t}$.
II.4 Correction to the soft waveform
A soft waveform is the exact zero-frequency limit of the spectrum in the presence of memory. In reality, we observe the spectrum at a finite frequency, and thus will see a deviation from the soft waveform. There are no generic constraints on such corrections other than that they vanish at $f=0$, since an arbitrary waveform with zero memory can be superposed on the original waveform without affecting the memory. The correction therefore contains information from the bulk of the time interval, and the leading-order (LO) correction encodes the next-to-leading order term in the asymptotic time-domain waveform.
We focus on the case of displacement memory, and define the correction factor to be
$\mathcal{C}(f)\equiv\left(\frac{i\Delta h}{2\pi f}\right)^{-1}\tilde{h}_{\infty}(f)-1.$ (14)
For all three models in Sec. II.1, $\mathcal{C}(f)$ is analytic at $f=0$. This is generally not the case for realistic waveforms. E.g., $\lim_{t\to\pm\infty}h(t)\to-1/(2|t|)$ corresponds to $\tilde{h}_{\infty}(f)=\ln|f|$. Due to the finite signal duration and instrumental noise (the latter usually being the more limiting factor), the spectrum at $f<f_{\text{min}}$ is inaccessible. It is thus more convenient to write a generic spectrum as
$\tilde{h}_{\infty}\equiv\frac{iHe^{-2i\varsigma}}{2\pi f}[1+C(f)],$ (15)
with $H>0$, $\varsigma\in\mathbb{R}$, and $C(f_{\text{min}})=0$. As $f_{\text{min}}\to 0$ (which requires the measurement time $T\to\infty$), $\{\varsigma,H\}\to\{0,\Delta h\}$, if the memory truly persists. Other reasonable definitions of $H$ are possible, but we are interested in situations where $\Delta h$ is not too different from $i2\pi f_{\text{min}}\tilde{h}_{\infty}(f_{\text{min}})$ (otherwise the soft waveform would be a poor approximation) so the corresponding correction factors $C(f)\equiv\left(\frac{iH}{2\pi f}\right)^{-1}\tilde{h}_{\infty}(f)-1$ should not differ significantly. Note that $C(f)$ cannot generally be obtained from the Taylor expansion about $f_{\text{min}}$, which may not converge away from $f_{\text{min}}$. If both linear polarization components are described by soft waveforms of the same type, and their arrival times $t_{*}^{\lambda}$ satisfies $f|(t_{*}^{+}-t_{*}^{\times})|\ll 1$, then the GW is approximately linearly polarized (i.e., $\tilde{h}_{+}/\tilde{h}_{\times}\in\mathbb{R}$). This may not hold in general due to finite-frequency corrections. Nevertheless, we define $\tilde{h}_{+}-i\tilde{h}_{\times}=\frac{iH_{+}e^{-2i\varsigma_{+}}}{2\pi f}[1+C_{+}(f)]+\frac{H_{\times}e^{-2i\varsigma_{\times}}}{2\pi f}[1+C_{\times}(f)]\equiv\frac{iHe^{-2i\psi}}{2\pi f}[1+C(f)]$, where $H$ and $C(f)$ are the effective memory amplitude and correction factor, respectively.
III Memory of compact binaries
In this section, we examine examples of soft waveforms of displacement memory produced by compact binaries. Before considering the concrete waveforms, we briefly discuss the effects of GW propagation on the observed soft waveforms in the eikonal approximation.
In an asymptotically flat spacetime, the gravitational waveform of a point source in the limit to null infinity can be described in the radiative coordinates as
$h(u=t-R)=h_{+}-ih_{\times}\equiv\sum_{l\geq 2,m}h_{l,m}(u)\,_{-2}Y_{lm}(\Theta,\Phi),$ (16)
with $R$ being the observer’s distance to the source, $(\Theta,\Phi)$ the spherical polar coordinates of the observer, and $_{s}Y_{lm}$ the spin-weighted spherical harmonics with spin weight $s$. Here the TT polarization is defined by $\mathbf{a}=\mathbf{e}_{\Theta}$ and $\mathbf{b}=\mathbf{e}_{\Phi}$ in Eq. (3). We also use the Cartesian coordinates $(X,Y,Z)=R(s_{\Theta}c_{\Phi},s_{\Theta}s_{\Phi},c_{\Theta})$, where $c_{z}\equiv\cos z$ and $s_{z}\equiv\sin z$.
The waveform modes $h_{l,m}(u)$ are expressed in terms of the radiative mass and current multipole moments $U_{lm}$ and $V_{lm}$ as $h_{l,m}=\left(U_{lm}-iV_{lm}\right)/(\sqrt{2}\,R)$ [30]. At leading post-Newtonian (PN) order, the $l=2$ modes correspond to the Newtonian-order (0PN) quadrupole waveform as given by the formula:
$h_{ij}=\frac{2}{R}\,\Lambda_{ij,kl}\,\ddot{M}_{kl},$ (17)
where $M_{ij}\equiv\int d^{3}x\,\rho\,x_{i}x_{j}$ is the Newtonian mass quadrupole moment tensor, and $\Lambda_{ij,kl}\equiv\frac{1}{2}(P_{ik}P_{jl}+P_{il}P_{jk}-P_{ij}P_{kl})$ is the TT projector, with $P_{ij}\equiv\delta_{ij}-\hat{k}_{i}\hat{k}_{j}$.
III.1 Propagation of waveforms
Eq. (16) is an idealized description, since in reality, the spacetime is not asymptotically flat. For a locally plane GW with sufficiently high frequency, [^1] Here we consider a frequency band that is low enough for the soft waveform to be a good approximation, while still high enough for the eikonal approximation to be applicable. its propagation in a curved spacetime background can be described by the eikonal approximation. If the observed frequency of a Fourier mode with original frequency $f$ is shifted to $f^{\prime}=\varrho f$, the observed frequency spectrum $\tilde{h}_{\infty}^{\prime}(f)$ is related to the original spectrum $\tilde{h}_{\infty}(f)$ via
$\tilde{h}_{\infty}^{\prime}(f)=(F/\varrho)\,\tilde{h}_{\infty}(f/\varrho),$ (18)
where $F$ is the amplitude modulation factor (which becomes frequency-dependent in the wave-optics regime [83, 12]).
The zeroth order eikonal approximation of massless relativistic waves due to a curved spacetime background is captured by the massless Klein-Gordon equation in the background coordinates. We consider a spatially flat Friedmann-Lemaitre-Robertson-Walker spacetime that is weakly perturbed: ${ds^{2}}/{a^{2}}=-(1+2\Psi)d\eta^{2}+2\Xi_{i}d\eta dx^{i}+(\delta_{ij}+H_{ij})dx^{i}dx^{j}$, where $a$ is the cosmological scale factor, $d\eta\equiv dt/a$ is the conformal time. The massless Klein-Gordon equation in the coordinates $(\eta,\mathbf{x})$ reads ${\partial_{a}\left(\sqrt{-g}\,g^{ab}\partial_{b}\phi\right)}/{\sqrt{-g}}=(W+B^{a}\partial_{a})\phi=0$, where $W\equiv g^{ab}\partial_{a}\partial_{b}$ and $B^{a}\equiv\partial_{b}(\sqrt{-g}\,g^{ab})/\sqrt{-g}$. In the eikonal approximation [90], the complex scalar field is described by a rapidly varying real phase $S(\eta,\mathbf{x})$ and a slowly varying real amplitude $A(\eta,\mathbf{x})$ as $\phi=Ae^{iS}$, with $\omega(\eta,\mathbf{x})\equiv-\partial_{\eta}S=2\pi f$ and $\mathbf{k}(\eta,\mathbf{x})\equiv\nabla S$. At zeroth order, $W\phi\approx 0$, and $k_{a}=\partial_{a}S=(-\omega,\mathbf{k})$ satisfies the local dispersion relation $D(x,k)\approx-g^{ab}k_{a}k_{b}\approx(1-2\Psi)\omega^{2}-|\mathbf{k}|^{2}+H_{ij}k_{i}k_{j}+2\Xi_{i}\omega k_{i}=0$. $k_{a}(\eta,\mathbf{x}(\eta))$ evolves according to ${dk_{a}}/{d\eta}\approx(\partial_{a}D)/(\partial_{\omega}D)$, e.g.,
$\frac{d\omega}{d\eta}\approx\frac{\omega^{2}\partial_{\eta}\Psi-\frac{1}{2}(\partial_{\eta}H_{ij})k_{i}k_{j}-(\partial_{\eta}\Xi_{i})\omega k_{i}}{\omega}.$ (19)
Meanwhile, the amplitude evolves according to $\left[(\partial_{k_{a}}D)\partial_{a}+\left(\partial_{k_{a}}\partial_{k_{b}}D\right)(\partial_{a}k_{b})/2-B^{a}k_{a}\right]A\approx 0$, i.e., $k^{a}\partial_{a}\ln A\approx-\nabla_{a}k^{a}/2$. The frequency measured by a timelike observer with four-velocity $u^{\mu}$ (normalized as $u^{\mu}u_{\mu}=-1$) is then $\omega_{\text{o}}=-u^{\mu}k_{\mu}$. E.g., for an observer (emitter) with $u_{\text{o(e)}}^{\mu}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\delta^{\mu}_{0}$ at the time $\eta_{\text{o(e)}}$ of observation (emission), $\omega_{\text{o(e)}}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\,\omega(\eta_{\text{o(e)}})$. In the absence of metric perturbations, $\omega(\eta_{\text{o}})=\omega(\eta_{\text{e}})$, thus $\omega_{\text{o}}=\omega_{\text{e}}/(1+z)$ with $z\equiv a_{\text{o}}/a_{\text{e}}-1$ being the redshift of the source.
For a spherical wave, the amplitude decay $\propto 1/(a_{\text{o}}R)$ needs to be included, where $R$ is the distance in the comoving coordinates $\mathbf{x}$, related to the source’s luminosity distance $d_{L}=z/H_{0}+\mathcal{O}(z^{2})$ via $a_{\text{o}}R=d_{L}/(1+z)$, $H_{0}$ being the Hubble parameter today. According to Eq. (18) with $F=1$, a waveform $\tilde{h}(f,R)$ in an asymptotically flat spacetime is converted to $\tilde{h}^{\prime}(f,d_{L})=(1+z)\tilde{h}((1+z)f,d_{L}/(1+z))=(1+z)^{2}\tilde{h}((1+z)f,d_{L})$. In the case of displacement memory, this implies that $d_{L}\Delta h^{\prime}(d_{L})=(1+z)R\Delta h^{\prime}(R)$, where $\Delta h^{\prime}(R)$ is observed at a distance $R$ sufficiently close to the source where the approximation of an asymptotically flat spacetime is good. This is consistent with the analysis of [6]. In summary, the effect of cosmic expansion on the soft waveform of displacement memory is a memory amplitude $\Delta h^{\prime}(d_{L})=(1+z)\Delta h(R=d_{L})$, where $\Delta h$ is computed in an asymptotically flat spacetime with a source distance $R=d_{L}$; meanwhile, as the whole spectrum shifts to lower frequencies, the deviation from the soft waveform at a given frequency is likely to increase.
For an observer (or a source) with non-vanishing velocity $u^{i}$, an additional apparent frequency change appears, which in a flat background is just the usual Doppler shift, given by $\varrho=(1-|\mathbf{v}|^{2})^{-1/2}(1-\hat{\mathbf{k}}\cdot\mathbf{v})$, with $\mathbf{v}=\mathbf{v}_{\text{o}}-\mathbf{v}_{\text{e}}$. As an example of the effect of metric perturbations, consider the wave traveling in $H_{ij}(u=\eta-\mathbf{v}_{g}\cdot\mathbf{x})$, where $\mathbf{v}_{g}$ is constant. Since $\frac{d}{d\eta}\approx\partial_{\eta}+\hat{k}_{i}\partial_{i}$ along the propagating direction $\hat{\mathbf{k}}$ of the wave, $\int_{\eta_{1}}^{\eta_{2}}d\eta\,\partial_{\eta}X(\eta,\mathbf{x})\approx-\frac{\Delta X}{1-\mathbf{v}_{g}\cdot\hat{\mathbf{k}}}$, where $\Delta X\equiv X(\eta_{2},\mathbf{x}_{2})-X(\eta_{1},\mathbf{x}_{1})$. From Eq. (19), the accumulated frequency shift is thus
$\varrho-1=\frac{\Delta\omega}{\omega}=\frac{1}{\omega}\int d\eta\frac{d\omega}{d\eta}\approx-\frac{\Delta\left(\frac{1}{2}\hat{k}_{i}\hat{k}_{j}H_{ij}\right)}{1-\mathbf{v}_{g}\cdot\hat{\mathbf{k}}}$. (20)
III.2 Displacement memory from compact binary scattering
The displacement and velocity memory can appear at 0PN in a linear gravitational radiation, since according to the quadrupole formula (17), they can be produced respectively by the asymptotic changes of the second and third time derivatives of $M_{ij}$:
$\ddot{M}_{ij}$ $=\int d^{3}x\,\rho\,(2v_{i}v_{j}+a_{i}x_{j}+x_{i}a_{j}),$ (21)
$\dddot{M}_{ij}$ $=\int d^{3}x\,\rho\,(3a_{i}v_{j}+3v_{i}a_{j}+\dot{a}_{i}x_{j}+x_{i}\dot{a}_{j}),$
where $v_{i}={dx^{i}}/{dt}$ is the velocity and $a_{i}={dv^{i}}/{dt}$ is the acceleration of the nonrelativistic matter. Velocity memory can thus be produced by an asymptotic change of acceleration accompanied by nearly constant velocity over the relevant time scale [43, 9], whereas displacement memory is produced in generic scattering processes. [^2] Similarly, the scattering of scalar or vector charges can also produce memory in the corresponding radiation (see Appendix B for a brief discussion). In this section, we examine the soft waveforms of compact binary scattering events in the moderately relativistic regime, where the PN approximation is applicable. We describe the scattering in the two-body center-of-mass (CM) frame, and denote $x_{I}^{\mu}=(t,\mathbf{x}_{I}(t))$ the 4-position of the $I$-th body with mass $m_{I}$. We also define $\mathbf{x}\equiv\mathbf{x}_{2}-\mathbf{x}_{1}$, $\mathbf{v}\equiv\dot{\mathbf{x}}$ and $r=|\mathbf{x}|$. In the CM frame, the initial state is fully specified by the relative velocity $\mathbf{v}_{0}=\mathbf{v}(t\to-\infty)$ and the impact parameter vector $\mathbf{b}=[\mathbf{r}-(\mathbf{r}\cdot\hat{\mathbf{v}})\,\hat{\mathbf{v}}]_{t\to-\infty}$, with $\hat{\mathbf{v}}=\mathbf{v}/|\mathbf{v}|$; we denote $v_{0}=|\mathbf{v}_{0}|$ and $b=|\mathbf{b}|$. Covariantly, the initial state can be defined by the four-velocities $u_{I}^{\mu}=(1-|\mathbf{v}_{I}|^{2})^{-1/2}(1,\mathbf{v}_{I})$ with an impact parameter four-vector $b^{\mu}$ satisfying $b_{\mu}u_{1}^{\mu}=b_{\mu}u_{2}^{\mu}=0$. The CM frame is defined by $m_{1}u_{1}^{i}+m_{2}u_{2}^{i}=0$, in which $b^{\mu}=(0,\mathbf{b})$.
We consider first the quadrupole waveform of a Newtonian hyperbolic encounter. The orbit of $\mathbf{x}(t)$ is taken to be a Keplerian orbit with semi-major axis $a$ and eccentricity $e$. Correspondingly, $b=a\sqrt{e^{2}-1}$, $v_{0}=\sqrt{M/a}$; so $e=\sqrt{1+(b/M)^{2}v_{0}^{4}}$, $a=M/v_{0}^{2}$. This approximation is reasonable if the incident velocity $v_{0}$ (for a given impact parameter $b$) is neither too large nor too small, since in the former case the conservative PN corrections become important, while in the latter case the gravitational radiation reaction matters. [^3] For example, a Keplerian parabolic orbit produces only null displacement memory [32]. However, when gravitational radiation reaction is included, a final state with vanishing relative velocity arises from nonzero initial relative velocity, leading to nonzero ordinary displacement memory. The quadrupole moment in the CM frame is given by $M_{ij}=\nu Mx_{i}x_{j}$, where $M=m_{1}+m_{2}$ is the total mass and $\nu=m_{1}m_{2}/M^{2}\in(0,1/4]$ is the symmetric mass ratio.
We choose a source frame in which the orbital angular momentum aligns with the $Z$-axis, with the periastron on the $X$ axis; we also shift the time coordinate such that $t=0$ corresponds to the periastron crossing. With this choice, only $h_{2,\pm 2}$ and $h_{2,0}$ are non-vanishing. Using the eccentric-anomaly parameterization of the hyperbolic orbit (see, e.g., [15]) and the integral representation of the Hankel function of the first kind: $H_{m}^{(1)}(z)=\frac{1}{i\pi}\int_{-\infty}^{\infty}d\xi\,e^{z\sinh\xi-m\xi}$, the idealized spectrum of the quadrupole waveform can be derived as
$\frac{f\tilde{h}_{2,\pm 2}/\varpi}{2\pi h_{0}}$ $=\left(\frac{\pm\sqrt{e^{2}-1}}{2e}+\frac{e^{2}-1}{e}\varpi\right)H_{1+in}$ (22)
$\quad+\left[\left(\frac{i\mp\sqrt{e^{2}-1}}{2e^{2}}-\frac{i}{4}\right)\right$.
$\left.\quad-\frac{e^{2}-1}{e^{2}}\left(1\pm i\,\sqrt{e^{2}-1}\right)\varpi\right]H_{in},$
$\frac{f\tilde{h}_{2,0}}{h_{0}}=-\frac{i\pi}{\sqrt{6}}\varpi H_{in},$ (23)
with $h_{0}\equiv\sqrt{{64}/{(5\pi)}}\,{\nu M^{2}}/{(aR)}$, $\varpi\equiv\pi f/\Omega$, $\Omega\equiv\sqrt{M/a^{3}}$, $n\equiv 2\varpi$, and $H_{m}\equiv H_{m}^{(1)}(ine)$. The relevant spin-weighted spherical harmonics are
$\left(\begin{smallmatrix}{}_{-2}Y_{2,2}\\ {}_{-2}Y_{2,0}\\ {}_{-2}Y_{2,-2}\end{smallmatrix}\right)=\sqrt{\frac{5}{4\pi}}\left(\begin{smallmatrix}e^{2i\Phi}c^{4}_{\Theta/2}\\ \sqrt{\frac{3}{8}}s_{\Theta}^{2}\\ e^{-2i\Phi}s^{4}_{\Theta/2}\end{smallmatrix}\right).$ (24)
Using the limits: $\lim_{n\to 0}H_{in}\to i\left(\frac{2}{\pi}\ln ne\right)$ and $\lim_{n\to 0}H_{1+in}\to\left(-\frac{2}{\pi en}-\frac{1}{e}\right)+i\left(\frac{2}{\pi e}\ln ne\right)$, we obtain from Eqs. (22)-(23) the soft waveform $\tilde{h}(f)\to{i(\Delta h_{+}-i\Delta h_{\times})}/{(2\pi f)}$ (since the periastron crossing happens at $t=t_{*}=0$) with its LO correction:
$\frac{f\tilde{h}_{2,\pm 2}}{h_{0}}\approx$ $\mp\frac{\sqrt{e^{2}-1}}{e^{2}}(1+\pi\varpi)+2\left(\frac{1}{e^{2}}-1\right)\varpi$ (25)
$+\left(1-\frac{2}{e^{2}}\right)\varpi\,\ln 2e\varpi,$
$\frac{f\tilde{h}_{2,0}}{h_{0}}\approx$ $\sqrt{\frac{2}{3}}\,\varpi\ln 2e\varpi.$
The correction factor depends on $\varpi$ and $e$. It turns out that the LO correction to $\text{Im}\,\tilde{h}_{\lambda}$ is linear, with $a_{0}=\pi^{2}/\Omega$; while the LO correction to $\text{Re}\,\tilde{h}_{\lambda}$ is logarithmic. These corrections are associated with the $1/t$ decay of the waveform as $t\to\pm\infty$ (see Sec. II). As a measure of the relative deviation from the soft waveform, we show the contours of $|f\tilde{h}_{2,\pm 2}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})-1|$ and $|f\tilde{h}_{2,0}/(\lim_{f\to 0}f\tilde{h}_{2,\pm 2})|$ in Fig. 1. The deviations decrease with $e$ in the shown parameter space, and become less than 0.1 at $f\lesssim 10^{-2}\Omega/\pi$. The deviation of $\text{arg}\,\tilde{h}$ at given $\{\Theta,\Phi\}$ from the soft limit is smaller (see Fig. 2 for a concrete example). Therefore, a soft waveform with a real correction factor $C(f)$ provides a good approximation to $\tilde{h}(f)$ at sufficiently low frequencies.
From Eq. (25), the memory amplitude $|\Delta h|=\sqrt{\Delta h_{+}^{2}+\Delta h_{\times}^{2}}$ is given by
$\frac{|\Delta h|}{\frac{\nu M}{R}\frac{\sqrt{e^{2}-1}}{ae^{2}}}=\sqrt{-8s^{4}_{\Theta}c_{4\Phi}+28c_{2\Theta}+c_{4\Theta}+35},$ (26)
thus $|\Delta h|$ is proportional to the quantity:
$\frac{M\sqrt{e^{2}-1}}{ae^{2}}=\frac{b/M}{(b/M)^{2}+1/v_{0}^{4}},$ (27)
which increases with $v_{0}$ for given $b$, and peaks at $b=M/v_{0}^{2}$ for given $v_{0}$. For given $b$ and $v_{0}$, $|\Delta h|$ takes the maximum value $8$ at $\Theta\in\{0,\pi\}$, and vanishes at $\Theta=\pi/2\,\land\,\Phi\,\text{mod}\,\pi/2=0$.
In the high-frequency regime, a “hard” waveform can be derived from Eqs. (22)-(23) by using the limits [37]:
$\lim_{n\to\infty}H_{in}$ $\to-\frac{ie^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{(e^{2}-1)^{1/4}\sqrt{n\pi/2}},$ (28)
$\lim_{n\to\infty}H_{1+in}$ $\to-\frac{\left(\sqrt{e^{2}-1}+i\right)e^{-n\left(\sqrt{e^{2}-1}-\text{arcsec}\,e\right)}}{e(e^{2}-1)^{1/4}\sqrt{n\pi/2}}.$
Figure 1: Relative deviations from the soft waveform in $\tilde{h}_{2,2}$ (black), $\tilde{h}_{2,-2}$ (blue) and $\tilde{h}_{2,0}$ (green).
Figure 2: Relative deviations from the soft waveform in $\text{arg}\,\tilde{h}$ for various eccentricities, at $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$.
Many of the features outlined above persist once PN corrections are included, at least when those corrections are not too large. We include the PN corrections to the CM-frame equation of motion (EOM) $\ddot{\mathbf{x}}=-(M/r^{3})\,\mathbf{x}+\mathbf{F}$ and the gravitational waveform according to [60]. The initial condition is set by an osculating hyperbolic orbit with orbital elements $\{a,e\}$ and the total mass parameter $M$ (see, e.g., [15]). Such a choice for the osculating orbit is not unique, since the initial orbit depends only on $\mathbf{v}_{0}$ and $\mathbf{b}$. However, this setting enable us to directly compare the results of Newtonian and PN EOMs in the harmonic coordinates. We choose the initial conditions such that the osculating Keplerian orbit coincides with the one considered above. For simplicity, we neglect the spins of the two bodies, thus $\mathbf{F}=\mathbf{F}_{\text{1PN}}+\mathbf{F}_{\text{2PN}}+\mathbf{F}_{\text{2.5PN}}+\mathcal{O}(\text{3PN})$; the PN orbit also lies in the $XY$ plane, but its periastron deviates from that of the Keplerian orbit. At 1PN order, the change of the argument of periastron $\varphi_{0}$ of the osculating orbit during the scattering is given by [15, 11]
$\frac{\Delta\varphi_{0}}{M/a}=\frac{(5\nu+2)e^{2}-5\nu+4+\frac{12e^{2}\arctan\frac{e+1}{\sqrt{e^{2}-1}}}{\sqrt{e^{2}-1}}}{e^{2}\sqrt{e^{2}-1}}.$ (29)
For concreteness, we choose $\nu=0.2$, $b/M=100$ and $v_{0}=0.13$, the results are shown in Fig. 3. To verify the convergence of PN corrections, we perform the computation at different PN orders. The approximation (29) (shown as the thick gray line in the top panel of Fig. 3) is found to agree well with the 2.5PN orbit, indicating that higher-order PN corrections are relatively small. The orbit is therefore well described by the 1PN-accurate quasi-Keplerian parameterization [23]. Nevertheless, there is a considerable difference between the 0PN and 0.5PN waveforms, but the 1PN correction to the waveform is already small. It should be noted that the null memory contribution to the waveform in the present case enters at 2.5PN order [32], and is thus negligible.
As a test, we verify that the spectrum of 0PN waveform computed using 0PN orbit agrees perfectly with Eqs. (22)-(23). The spectra of $|\text{Re}\,\tilde{h}_{\pm}(f)|$ and $|\text{Im}\,\tilde{h}_{\pm}(f)|$ of the 1PN waveforms are shown in the bottom panel of Fig. 3. The numeric spectrum at the lowest frequency is pretty close to the soft waveform with $\Delta h_{\lambda}$ computed in the time domain. The spectrum of $f\,\text{Im}\,\tilde{h}_{\pm}$ at low frequencies is well described by a linear correction: $f\,\text{Im}\,\tilde{h}_{\lambda}\propto 1+\kappa_{\lambda}\varpi$, with $\kappa_{+}\approx 0.95\pi$ and $\kappa_{\times}\approx 0.92\pi$. These slightly deviate from the Keplerian value $\kappa=\pi$ when expanded near $f=0$, as given by Eq. (25). Notably, the correction factors of $f\,\text{Im}\,\tilde{h}_{+}(f)$ and $f\,\text{Im}\,\tilde{h}_{\times}(f)$ are very close to each other at all frequencies. As the frequency increases, the real part of the spectrum becomes more important.
At sufficiently high frequency, $\text{Im}\,\tilde{h}_{+}\approx\text{Re}\,\tilde{h}_{\times}$ and $\text{Re}\,\tilde{h}_{+}\approx\text{Im}\,\tilde{h}_{\times}$, indicating that the radiation is almost purely right-handed polarized. This is also reflected in the 0PN spectra of the $Z$-component angular momentum and graviton number fluxes, whose ratio approaches 2 at high frequencies, as for an elliptical orbit. The decaying behavior at high frequencies is similar to the 0PN approximation given by Eq. (28), as depicted by the green lines in the bottom panel of Fig. 3.
Fig. 4 shows the spectrum $\tilde{h}^{\prime}(f)$ obtained by an overall phase rotation that transforms $\tilde{h}$ at the lowest frequency of the numeric spectrum to a purely imaginary value. We see that $\tilde{h}^{\prime}(f)$ is well described by a soft waveform with a real correction factor $C(f)=1+\kappa\varpi+\mathcal{O}(\varpi^{2})$ at low frequencies. In this example, $\kappa\approx 2.63\pi$, which is quite different from $\kappa_{\lambda}$ describing the linear correction to $f\,\text{Im}\,\tilde{h}_{\lambda}$.
Figure 3: CM-frame orbits (top panel) and gravitational waveforms in the time domain (middle panel) and frequency domain (bottom panel) of a compact binary scattering, with $\nu=0.2$, $b/M=100$, $v_{0}=0.13$ (corresponding to $e=2$), and $\{\Theta,\Phi\}=\{\pi/6,\pi/3\}$. In the middle panel, the 0PN waveform is computed using the 0PN orbit, whereas the 0.5PN and 1PN waveforms are computed using the 2PN orbit. The bottom panel shows the spectrum of the 1PN waveform, with $\Omega=\sqrt{M/a^{3}}=v_{0}^{3}/M\approx\pi(v_{0}/0.13)^{3}(100\,M_{\odot}/M)\times 1.4\,\text{Hz}$. Before computing the spectrum, we shift the time coordinate such that the periastron crossing happens at $t=0$. The horizontal line in the bottom panel represents the soft waveform corresponding to $\Delta h_{\lambda}$ in the time domain.
Figure 4: Spectra of $f\,\text{Im}\,\tilde{h}^{\prime}(f)$ and $f\,\text{Re}\,\tilde{h}^{\prime}(f)$. The linear and quadratic fits to the low-frequency part of the spectra are also shown. The horizontal line represents the soft waveform corresponding to $\Delta h_{\lambda}$ in the time domain.
For an extreme mass-ratio scattering involving a large black hole (BH), the orbit of the small compact body can be approximated by an unbound geodesic in the background BH spacetime. The dissipative effect is expected to be negligible in this case, since the incident velocity for scattering is bounded from below by the critical velocity for plunge in the test-body approximation. We compute the ordinary displacement memory produced by such a system using the general formula [10]:
$\Delta h_{ij}=\Lambda_{ij,kl}\sum_{I}\frac{4m_{I}}{R}\Delta\left[\frac{v_{I}^{k}v_{I}^{l}}{\sqrt{1-v_{I}^{2}}\,(1-\mathbf{v}_{I}\cdot\hat{\mathbf{k}})}\right$]. (30)
This is derived from the linearized Einstein equation in a flat background, but can be applied to scattering processes provided that the initial and final states live in a flat background. Eq. (30) can also be derived from the soft graviton theorem [78]. As can be checked, the low-velocity expansion of Eq. (30) up to $\mathcal{O}(v^{4})$ agrees with the 1PN approximation:
[…]
$\quad+(1-3\nu)\left[(\mathbf{v}\cdot\hat{\mathbf{k}})^{2}+v^{2}/2\right]\Big)\Big]\frac{2\nu M}{R},$
for $m_{2}\ll m_{1}$. For simplicity, we consider a non-spinning central BH, and thus compute the gauge-invariant deflection angle $\chi$ of the small test body with mass $\nu M$ in a Schwarzschild spacetime with mass parameter $M$. The source frame is again chosen such that the initial orbit lies in the $XY$ plane, the initial and final three-velocities are along the direction $\mathbf{N}_{\text{in}}=c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$ and $\mathbf{N}_{\text{out}}=-c_{\vartheta}\mathbf{e}_{X}+s_{\vartheta}\mathbf{e}_{Y}$, respectively, with $\vartheta=(\chi-\pi)/2$. The displacement memory from Eq. (30) is given by
$\frac{\Delta h_{+}}{F/R}$ $=\frac{c_{\chi+2\Phi}+v_{0}^{-2}}{1-v_{0}s_{\Theta}s_{\chi/2+\Phi}}-\frac{c_{\chi-2\Phi}+v_{0}^{-2}}{1+v_{0}s_{\Theta}s_{\chi/2-\Phi}}-\frac{s_{\Theta}s_{\chi/2}c_{\Phi}}{v_{0}/2},$ (32)
$\frac{\Delta h_{\times}}{F/R}$ $=\frac{v_{0}s_{2\Theta}s_{\chi/2}s_{\Phi}(c_{\chi}-c_{2\Phi})+2c_{\Theta}s_{\chi}c_{2\Phi}}{\left(v_{0}s_{\Theta}s_{\chi/2-\Phi}+1\right)\left(v_{0}s_{\Theta}s_{\chi/2+\Phi}-1\right)},$
where $F=2\omega$ and $v_{0}=1$ for a null particle with energy $\omega$, and $F=\frac{2mv_{0}^{2}}{\sqrt{1-v_{0}^{2}}}$ for a timelike particle with mass $m$. For $v_{0}<1$, Eq. (32) cannot be easily expanded in $_{-2}Y_{lm}$.
Figure 5: Ordinary displacement memory from the scattering of a test body with mass $\nu M$ by a non-spinning BH with mass $M$, for $b/M=100$, and $\{\Theta,\Phi\}=\{\pi/5,\pi/7\}$. The blue and black solid lines show the test-body approximation. The red (green) dashed lines show the 0PN (1PN) approximations. The pink dashed lines show the results computed from Eq. (32), using the 1PN approximation to the deflection angle. The critical incident velocity $v_{\text{plunge}}\approx 0.04$. The figure inset shows the results for small incident velocities, with the horizontal axis given by $v_{0}-v_{\text{plunge}}$. For comparison, the gray dot-dashed lines show the results for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, for which $v_{\text{plunge}}=0.02$.
The deflection angle depends on the initial velocity $v_{0}$ and the dimensionless impact parameter $b/M$. For the small body to be scattered rather than plunge into the BH, the impact parameter needs to exceed a critical value: ${b_{\text{plunge}}}/{M}=\sqrt{4+\frac{\sqrt{8v_{0}^{2}+1}-1}{2v_{0}^{4}}+\frac{4\sqrt{8v_{0}^{2}+1}+10}{v_{0}^{2}}}$. For a null particle, $b_{\text{plunge}}(v_{0}=1)=3\sqrt{3}M$. Likewise for a given impact parameter, $v_{0}$ needs to exceed a critical value $v_{\text{plunge}}$ in order to escape from the BH. It is possible for high-energy massless or ultra-relativistic particles (such as photons and neutrinos) to be scattered by a BH, but this cannot produce an observable memory, since even a photon with $\omega=10\,\text{TeV}$ amounts to a mass of only $9\times 10^{-54}\,M_{\odot}$.
For the scattering of timelike particles, a concrete example is shown in Fig. 5, where the results are also compared with the 0PN approximation in Eq. (25) (red dashed lines) and the results computed using Eq. (31) and the 1PN approximation to the deflection angle: $\chi=2\arccos(-1/e)-\pi+\Delta\varphi_{0}$ (green dashed lines), with $\Delta\varphi_{0}$ given by Eq. (29). For comparison, we also show the results (pink dashed lines) computed from Eq. (32), but using the 1PN approximation to the deflection angle, which remains in good agreement with the test-body approximation even at high velocities. As can be seen, the 0PN approximation is accurate only for intermediate values of $v_{0}$, and the situation near $v_{\text{plunge}}$ cannot be captured by the 1PN deflection angle either, where the produced memory is relatively small. For $v_{0}\gtrsim v_{\text{plunge}}$, the periastron occurs at $r\gtrsim 4M$; the result is therefore sensitive to the spin of the central BH. Eq. (32) can also be applied to the scattering orbit in the ecliptic plane of Kerr spacetime; the gray dot-dashed lines in Fig. 5 show the result for a central BH with dimensionless spin $\chi=1$ along the $Z$-axis, which are, interestingly, closer to the pink dashed lines. At high velocities, the amplitudes of memory are considerably larger than the 0PN prediction. As in the 0PN approximation, $|\Delta h|$ is invariant under the transformation $\Theta\to\pi-\Theta$ or $\Phi\to\pi-\Phi$, but its extreme point deviates from $\Theta\in\{0,\pi\}$. For $\Theta=\pi/2$, $|\Delta h|$ still vanishes at four values of $\Phi$, two of which occur at $\pi/2$ and $3\pi/2$, while the remaining two are displaced from 0 and $\pi$.
III.3 Displacement memory from quasi-circular non-precessing BBH merger
Due to the non-linearity of gravity, displacement memory can also be produced from the gravitational radiation itself [18]. Perturbatively, the spherical-harmonic modes of the produced secondary GW (memory modes) is given by the time integral [85, 30]: [^4] This can be obtained, e.g., from Eq. (30) by identifying the outgoing particles with a flux of gravitons [85], for which $\mathbf{v}_{I}=\mathbf{n}(\theta,\phi)$, $|\mathbf{n}|=1$, ${m_{I}}/{\sqrt{1-v_{I}^{2}}}\to\int dt^{\prime}d\Omega_{2}(\theta,\phi)\frac{dE_{\text{gw}}}{dt^{\prime}\,d\Omega_{2}}$, and using $\int d\Omega_{2}(\Theta,\Phi)\,\frac{\frac{1}{2}(e_{ij}^{+}-ie_{ij}^{\times})n_{i}n_{j}}{1-\mathbf{n}\cdot\hat{\mathbf{k}}(\Theta,\Phi)}\,_{-2}Y_{lm}^{*}(\Theta,\Phi)=4\pi\sqrt{\frac{(l-2)!}{(l+2)!}}Y_{lm}^{*}(\theta,\phi)$, where $e_{ij}^{\lambda}(\Theta,\Phi)$ is defined as in Eq. (16).
$h_{l,m}^{\text{mem}}(u)\approx\frac{16\pi}{R}\sqrt{\frac{(l-2)!}{(l+2)!}}\int^{u}dt^{\prime}\int d\Omega_{2}\,\frac{dE_{\text{gw}}}{dt^{\prime}\,d\Omega_{2}}\,Y_{lm}^{*}.$ (33)
Similar result also applies to other forms of null radiation. The angular energy flux distribution $\frac{dE_{\text{gw}}}{dt\,d\Omega_{2}}$ of the primary GW is determined by its energy density $\rho_{\text{gw}}=\left\langle\frac{1}{16\pi}|\dot{h}|^{2}\right\rangle$ in the wave zone, which can be written as a summation over $g=\{l_{1},m_{2},l_{2},m_{2}\}$ as [30]
$\frac{\frac{dE_{\text{gw}}}{dt\,d\Omega_{2}}}{R^{2}}=\rho_{\text{gw}}=\frac{1}{16\pi}\sum_{g}\left\langle\dot{h}_{l_{1},m_{1}}\dot{h}^{*}_{l_{2},m_{2}}\right\rangle\,_{-2}Y_{l_{1}m_{1}}\,{}_{-2}Y^{*}_{l_{2}m_{2}},$
where the angle brackets denote an average over several GW wavelengths. For a quasi-periodic GW source, the null memory thus grows linearly in time. As discussed in Sec. II.1, the detectability of such a waveform is poor. For the search of null displacement memory, one is therefore interested in the system capable of producing a strong and predictable GW burst, but with negligible ordinary displacement memory. The ideal systems to observe are the binary black hole (BBH) mergers. In this section, we discuss the potentially observable soft waveforms of such systems. Our focus is not on identifying their nonlinear nature (which is impossible without observing the primary GW), but rather on some features of their low-frequency spectra. We consider quasi-circular non-precessing BBH systems and employ the numerical relativity surrogate model NRHybSur3dq8_CCE [96] (constructed for the mass ratio $q=m_{2}/m_{1}\in[1,10]$) to compute their gravitational waveforms.
Long before the merger, 0PN approximation to the waveform is adequate. In this regime, the orbital evolution is given by $\dot{x}=Jx^{5}$, where $J\equiv 64\nu/(5M)$, the PN parameter $x\equiv(M\Omega)^{2/3}$, with $\Omega$ being the orbital frequency. Meanwhile, the gravitational radiation in a source frame with $Z$-axis aligned with the orbital angular momentum is dominated by the $h_{2,\pm 2}$ modes, corresponding to the quadrupole waveform. Due to the adiabaticity of orbital evolution, the memory mode after its steady accumulation also appears at 0PN order, and is dominated by the $h_{2,0}$ mode. This leads to the LO memory-mode waveform [92] with $h_{l,m}^{\text{(mem)}}(x=0)=0$: $h_{+}^{\text{(mem)}}(t)=Kx$, $h_{\times}^{\text{(mem)}}(t)=0$, where $K\equiv{\nu Ms^{2}_{\Theta}\,(17+c^{2}_{\Theta})}/(48R)$. At $t\geq t_{0}$, we are only interested in the change of $h_{+}^{\text{(mem)}}$ with respect to its initial value at $t_{0}$, which is given by $\Delta h_{+}^{\text{(mem)}}(t)\equiv h_{+}^{\text{(mem)}}(t)-h_{+}^{\text{(mem)}}(t_{0})=Kx_{0}\left[1-(1+4Jx_{0}^{4}\Delta t)^{-1/4}\right]$, where $x_{0}=x(t=t_{0})$ and $\Delta t=t-t_{0}$. A concrete example is presented in Fig. 6. As shown in the bottom panel, the initial memory-mode waveform is well-described by this 0PN approximation, which in turn can be approximated by a linear growth (expected for a quasi-periodic GW source) initially. The Newtonian regime of inspiral is followed by the late inspiral, merger, and ringdown, during which the memory accumulates much faster than the 0PN prediction before finally settling to a saturated value. Throughout this process, $h_{2,0}$ is the dominating memory mode. The top panel of Fig. 6 compares the memory-mode waveform with the full (memory + oscillatory) waveform. Although the oscillatory waveform is the absolutely dominating component before the final merger, the memory-mode waveform is the only final remenant of this event, if the observation starts before the merger. This is manifested in the frequency spectrum of the full waveform shown in Fig. 7, whose low-frequency part is contributed solely by the memory modes. In contrast, the low-frequency spectrum of $h_{\times}(t)$, dominated by the spin memory mode $h_{30}$ [69], is much weaker than that of $h_{+}(t)$. In practice, the low-frequency signal from spin memory would therefore be masked by that of displacement memory.
As shown in the upper figure inset in Fig. 7, $\text{Im}\,\tilde{h}_{+}$ of the memory-mode waveform is almost purely imaginary until the frequency is sufficiently high, at which the spectrum has been overwhelmed by the oscillatory component that appears after $f\gtrsim\Omega(t_{0})/\pi$. Thus, a soft waveform with a negative real correction factor $C(f)$ provides a good approximation to $\tilde{h}(f)$ at low frequencies.
Figure 6: Gravitational waveform of a quasi-circular BBH merger from NRHybSur3dq8_CCE, with $q=1$, $M=100\,M_{\odot}$, $\Theta=\pi/3$ and $R=400\,\text{Mpc}$ (the effect of cosmic expansion is not included). The top panel compares the full waveform (orange) with the memory-mode waveform due to $h_{2,0}$ and $h_{40}$ modes (purple), and the 0PN approximation to the oscillatory waveform (black dashed), for a non-spinning system. The black dot-dashed line shows the memory-mode waveform for the case of aligned dimensionless spins $\chi_{1,2}=0.8$. The blue dot-dashed line shows the 0PN approximation to the memory-mode waveform for an elliptical orbit [32] with eccentricity $e(t_{0})=0.05$ and semi-major axis $a(t_{0})=M/x_{0}$. The lower panel compares the $h_{2,0}$ mode (blue dashed) with its 0PN approximation (red dashed) and linear approximation (black dashed), for a non-spinning system.
Figure 7: Frequency spectra of the waveforms in Fig. 6. The horizontal red (black) dashed line represents the soft waveform corresponding to $\Delta h_{+}$ of the full (reduced) time-domain waveform. The black solid line is the spectrum of the shortened waveform. We also show a local fit to the spectrum of the full waveform by a two-parameter logarithmic model $A+B\ln f$ (blue dashed line), the memory-mode spectrum for $q=8$ multiplied by 6.28 (green dot-dashed line), and the memory-mode spectrum for $q=1$ and aligned spins $\chi_{1,2}=0.8$ multiplied by 0.49 (black dot-dashed line). The blue dot-dashed line shows the 0PN oscillatory-mode spectrum ($f|\tilde{h}_{+}|\propto f^{-1/6}$) in the stationary phase approximation. The lower figure inset shows a zoomed-in view of the low-frequency region. The upper figure inset compares $f\,|\tilde{h}_{+}(f)|$ with $f\,|\text{Im}\,\tilde{h}_{+}(f)|$ (red dot-dashed line) of the memory-mode waveform.
It should be stressed that in this example the low-frequency spectrum does not faithfully reflect the memory amplitude, which is precisely due to the slow accumulation of memory during the early inspiral described above. The horizontal red dashed line in Fig. 7 indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})|$, while the black dashed line indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})-\Delta h_{+}^{\text{(mem)}}(t_{\text{mid}})|$, which is the memory height of the shortened waveform: $h_{+}^{\prime}(t)=[h_{+}(t)-h_{+}(t_{\text{mid}})]\,\Theta(t-t_{\text{mid}})$ $t_{\text{mid}}=t_{\text{merger}}-300\,\text{s}$ chosen so that $h_{2,0}$ can be described by the 0PN approximation for $t<t_{\text{mid}}$. We verified that the numerical spectrum of the waveform $h_{+}(t)-h_{+}^{\prime}(t)$, describing the growth prior to $t_{\text{mid}}$, agrees well with the analytical spectrum derived from the LO memory waveform, validating the accuracy of the FFT]. Neither horizontal line precisely matches the low-frequency limit of its corresponding spectrum, since $10^{-4}\,\text{Hz}$ is still finite. Even if $t_{\text{merger}}-t_{0}$ is fixed but we look at sufficiently low frequencies (which can be achieved by enlarging $t_{\text{final}}-t_{\text{merger}}$), we see a soft waveform with the correct height $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})|$, which roughly becomes $h_{+}^{\text{(mem)}}(t_{\text{final}})-h_{+}^{\text{(mem)}}(t_{\text{mid}})+Kx(t_{\text{mid}})$ for $t_{0}\to$-$\infty$.
A small difference at the lowest frequencies is, however, the only place where the early evolution manifests itself, as the spectra of the shortened and full waveforms are barely distinguishable for $f>10^{-3}\,\text{Hz}$. Near the lowest frequencies, the spectrum can be approximated by a linear fit, whereas at moderately high frequencies it can be reasonably described by a logarithmic fit (which is much smaller than the linear correction observed at low frequencies), as shown by the blue dashed line in Fig. 7. For a given total mass $M$, as the mass ratio $q>1$ increases, the effective memory height decreases, while the relative correction $|C(f)|$ increases. This is illustrated in Fig. 7 by comparing the normalized spectra of systems with $q=1$ and $q=8$. Figs. 6-7 also illustrate that the total memory offset is insensitive to the orbital eccentricity in the early stage [32], while it can be significantly affected by the BH spins.
Enlarging $t_{\text{final}}-t_{\text{merger}}$ does not change the memory-mode spectrum, while enlarging $t_{\text{merger}}-t_{0}$ only leads to a small difference at the lowest frequencies if $x_{0}$ is small (as can be seen by comparing the shortened and full waveforms in Fig. 7). The impact of observation time on the memory-mode spectrum is thus small, provided that most of the memory is accumulated during the observational time interval. On the other hand, the lowest frequency covered by the oscillatory-mode spectrum is $f\approx\Omega(t_{0})/\pi$. In the 0PN regime, we have ${\Omega(t_{0}-\Delta t)}/{\Omega(t_{0})}=\ell$, with
$\ell\approx\left[1+26\left(\frac{\Omega(t_{0})}{\pi\text{Hz}}\right)^{\frac{8}{3}}\left(\frac{M/M_{\odot}}{100}\right)^{\frac{5}{3}}\left(\frac{\Delta t/\text{yr}}{10^{-3}}\right)\left(\frac{\nu}{\frac{1}{4}}\right)\right]^{-\frac{3}{8}}.$
Assuming that $\Omega(t_{0})/\pi=1\,\text{Hz}$, for $M=100\,M_{\odot}$ and $\nu=q/(1+q)^{2}=1/4$, $\ell(\Delta t=10^{-4}\,\text{yr})\approx 0.62$, $\ell(\Delta t=10^{-3}\,\text{yr})\approx 0.3$, while $\ell(\Delta t=1\,\text{yr})\approx 0.02$. For smaller $M$, $\Omega(t_{0})/\pi=1\,\text{Hz}$ corresponds to an earlier stage of the evolution. Therefore for the quasi-circular stellar-mass BBH mergers and $\Delta t<10^{-3}\,\text{yr}$, typically $\Omega(t_{0}-\Delta t)/\pi\sim 1\,\text{Hz}$.
Due to the mass-scaling symmetry of the orbital dynamics and gravitational waveforms for fixed dimensionless spins and mass ratio, the spectra of observed memory-mode waveforms of sufficiently long duration in an expanding Universe (see Sec. III.1) exhibit the usual mass–redshift degeneracy:
$\tilde{h}^{\prime}(f,M^{\prime},d_{L})=\left[\frac{(1+z)M^{\prime}}{M}\right]^{2}\tilde{h}\left(\frac{(1+z)M^{\prime}}{M}f,M,d_{L}\right).$
III.4 Constraint on the source from the displacement-memory amplitude
From Eq. (33), the total offset of vacuum null displacement memory at future null infinity in an asymptotically flat spacetime can be expressed by $\Delta h=\frac{2\pi}{R}\bar{\eth}^{2}\mathfrak{D}^{-1}\frac{dE_{\text{gw}}}{d\Omega_{2}}$, where $\bar{\eth}$ is the spin-lowering operator, $\mathfrak{D}=D^{2}(D^{2}+2)/8$, and $D^{2}$ is the spherical Laplacian. Including the ordinary contribution, the result is [67]
$\Delta h=\frac{1}{2}\bar{\eth}^{2}\mathfrak{D}^{-1}\Delta Q,$ (34)
with $\Delta Q=\left(\Delta m+4\pi\frac{dE_{\text{gw}}}{d\Omega_{2}}\right)\frac{1}{R}$, where $\Delta m$ is the change of Bondi mass aspect. For $l\geq 2$, Eq. (34) reads
$(\Delta h)_{lm}=4\sqrt{\frac{(l-2)!}{(l+2)!}}(\Delta Q)_{lm}.$ (35)
where $J_{lm}$ denotes the expansion coefficient of a spin-$s$ field $J(\Theta,\Phi)$ in terms of $_{s}Y_{lm}(\Theta,\Phi)$. Here we note that a constraint on the source can be derived from $|\Delta h|$,
$\mathcal{Q}$ $\equiv\int d\Omega_{2}\,|\Delta Q|^{2}-\sum_{l=0,1}\sum_{m}|(\Delta Q)_{lm}|^{2}$ (36)
$=\sum_{l\geq 2}\sum_{m}|(\Delta Q)_{lm}|^{2}\geq\frac{|\Delta h|^{2}}{\mathcal{I}(\Theta,\Phi)},$
with $\mathcal{I}=\sum_{l\geq 2}^{\infty}\sum_{m}\frac{16(l-2)!}{(l+2)!}\left|{}_{-2}Y_{lm}\right|^{2}$. For $\Theta=0$, $_{-2}Y_{lm}(\Theta,\Phi)=e^{im\Phi}\sqrt{\frac{2l+1}{4\pi}}\delta_{m,2}$, $\mathcal{I}=4/(3\pi)$, thus $\mathcal{Q}\geq(3\pi/4)|\Delta h|^{2}$. Taking into account the cosmic expansion discussed in Sec. III.1, this bound becomes $\mathcal{Q}_{R\to d_{L}}\geq(3\pi/4)|\Delta h|^{2}/(1+z)^{2}$.
IV Detectability Analysis
IV.1 TDI response
In this section, we analyze the soft memory signals at several planned space-based GW detectors. A LISA-like detector consists of three spacecraft (SC) forming an approximately equilateral triangle constellation, as depicted in Fig. 8. Laser beams are exchanged between the SC to form multiple interferometers. Six laser beams connecting the local and distant test masses interfere with the local lasers, allowing measurement of the frequency modulation of light that carries information about passing GWs.
We denote the position of the $I$-th SC by $\mathbf{x}_{I}(t)$. Consider the light sent from the $s$-th SC at time $t_{s}$ and which is received by the $r$-th SC at time $t$, i.e., $\mathbf{x}_{r}(t)=\mathbf{x}_{s}(t_{s})+L_{rs}(t)\,\mathbf{n}_{rs}(t)$, with the unit vector $\mathbf{n}_{rs}(t)\equiv[\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t_{s})]/|\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t_{s})|$ along the light propagation and the effective armlength $L_{rs}(t)\equiv|\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t_{s})|$. Since the motion of SC during the light-travel time is negligible, we can make the approximation $\mathbf{n}_{rs}(t)\approx[\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t)]/|\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t)|$, $L_{rs}(t)\approx|\mathbf{x}_{r}(t)-\mathbf{x}_{s}(t)|$ and $t_{s}\approx t-L_{rs}(t)$.
The propagation of laser light between the SC can be described in the leading-order eikonal approximation, as presented in Sec. III.1. The light frequency $\omega_{s}$ at $\mathbf{x}_{s}(t_{s})$ is fixed by the laser source, while the light frequency $\omega_{r}$ measured locally at $\mathbf{x}_{r}(t)$ is affected by both the motion of $\mathbf{x}_{r}$ [see Eq. (79) in Appendix B] and the light propagation. The effect of GW on the one-way Doppler shift $y_{rs}(t)\equiv[\omega_{r}(t)-\omega_{s}]/\omega_{s}$ can be analyzed conveniently in the TT coordinates, with the relevant metric given by Eq. (1). For light propagating between two static test-masses, the relative frequency modulation comes solely from the GW-induced modification of the photon dispersion relation, which for a plane wave $h_{ij}(t,\mathbf{x})=h_{ij}(u=t-\hat{\mathbf{k}}\cdot\mathbf{x})$ is given by Eq. (20):
$y_{rs}(t)=\frac{1}{2}n_{rs}^{i}n_{rs}^{j}\frac{h_{ij}(t-L_{rs}-\hat{\mathbf{k}}\cdot\mathbf{x}_{s})-h_{ij}(t-\hat{\mathbf{k}}\cdot\mathbf{x}_{r})}{1-\hat{\mathbf{k}}\cdot\mathbf{n}_{rs}},$ (37)
with $\hat{\mathbf{k}}$ denoting the unit vector in the direction of GW propagation.
In reality, the laser frequency contains random fluctuations. Such laser frequency noise, together with the acceleration noise of test masses and the noise of optical metrology system, constitute the main instrumental noises limiting the detector’s sensitivity to GWs. The ground-based GW interferometers do not suffer from the laser frequency noise due to their equal armlengths. For the space-based detectors, the laser frequency noise would be dominant in the measurements and should be mitigated by making suitable time-delayed linear combinations of the six one-way data streams $y_{rs}(t)$. This post-processing procedure called time-delay interferometry (TDI) [86] synthesizes a virtual equal-arm interferometer and thus suppresses the laser frequency noise. For example, the Michelson $X$ combination in the first-generation TDI is
$X_{1}(t)=\left(y_{13}+D_{13}y_{31}+D_{131}y_{12}+D_{1312}y_{21}\right)-(3\leftrightarrow 2),$ (38)
where the time delay operators are defined by
$D_{ij}x(t)$ $\equiv x(t-L_{ij}),$ (39)
$D_{i_{1}\cdots i_{n}}x(t)$ $\equiv\prod_{k=1}^{n-1}D_{i_{k}i_{k+1}}x(t),$ (40)
and for the computation of GW signals, it suffices to make the following approximation:
$D_{i_{1}\cdots i_{n}}x(t)\approx x\left(t-\sum_{k=1}^{n-1}L_{i_{k}i_{k+1}}\right).$ (41)
Figure 8: Left: The detector configuration in the comoving and corotating (xyzxyz) frame centered at the guiding center. Spacecraft (SC) are indexed 1, 2, 3 clockwise when looking down at their solar panels. SC 3 lies on the $y$-axis, $\mathbf{n}_{21}$ points in the negative $x$-direction. For LISA, Taiji and BBO Stage I, the constellation rotates clockwise about the $z$-axis at an approximately constant angular velocity $2\pi/\text{yr}$ in the SSB frame, while the guiding center travels in a quasi-circular heliocentric orbit on the ecliptic plane with radius $R\approx 1\,\text{AU}$. For TianQin, the SC follow geocentric orbits with a period of about 3.64 days [51]. Right: Schematic of the considered guiding center orbits of LISA and Taiji in the SSB ($x^{\prime}y^{\prime}z^{\prime}$) frame. The $z$-axis of the detector frame undergoes an approximate precession about the $z^{\prime}$-axis with a constant angular velocity $2\pi/\text{yr}$ and a fixed inclination angle $\angle(\mathbf{e}_{z},\mathbf{e}_{z^{\prime}})=\pi/3$. The LISA detector trails the Earth by $\pi/9$, while the Taiji detector leads the Earth by $\pi/9$ [74].
The TDI response is further simplified in the static equal-armlengths approximation (SEA), in which we neglect the motion of SC and use $L_{rs}\approx L=\text{const}$. The static-detector approximation is reasonable for a short-term observation, and the equal-armlengths approximation can also be sufficiently accurate in such cases, since the armlengths vary only at the percent level. Given the time-domain waveform $h_{ij}(t)=h_{ij}(t,\mathbf{0})$, the frequency spectrum of the time-domain TDI signals (with a sufficiently high sampling frequency) can be easily derived in this approximation. Using Eq. (37), the result for (38) is
$\tilde{X}_{1}(f)$ $=\frac{1}{2}\left[n_{13}^{i}n_{13}^{j}\left(F_{13}+e^{2\pi ifL}F_{31}\right)-(3\leftrightarrow 2)\right]\tilde{h}_{ij},$ (42)
where
$F_{rs}\equiv\left(1-e^{4\pi ifL}\right)\frac{e^{2\pi ifL(1-\hat{\mathbf{k}}\cdot\mathbf{n}_{rs})}-1}{1-\hat{\mathbf{k}}\cdot\mathbf{n}_{rs}}e^{2\pi if\,\hat{\mathbf{k}}\cdot\mathbf{x}_{r}}.$ (43)
Note that in the low-frequency limit ($fL\ll 1$),
$\tilde{X}_{1}(f)\approx 8\pi^{2}L^{2}f^{2}\sum_{\lambda}\left(n_{13}^{i}n_{13}^{j}-n_{21}^{i}n_{21}^{j}\right)\tilde{h}_{ij}\,e^{2\pi if\,\hat{\mathbf{k}}\cdot\mathbf{X}},$ (44)
where $\mathbf{X}$ is the guiding center of the constellation.
The other two Michelson combinations $Y$ and $Z$ can be obtained from $X$ through cyclic permutations $(1,2,3)\to(2,3,1)$ and $(1,2,3)\to(3,1,2)$, respectively. For the signal analysis, we consider the optimal combinations $\{A,E,T\}$:
$\left(\begin{matrix}A\\ E\\ T\end{matrix}\right)=\left(\begin{matrix}-\frac{1}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\\ \frac{1}{\sqrt{6}}&-\frac{2}{\sqrt{6}}&\frac{1}{\sqrt{6}}\\ \frac{1}{\sqrt{3}}&\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{3}}\end{matrix}\right)\left(\begin{matrix}X\\ Y\\ Z\end{matrix}\right).$ (45)
for which the noise correlation matrix is diagonalized in the idealized case [70]. In the SEA, $\tilde{T}(f)$ is strongly suppressed for $fL\ll 1$ relative to Eq. (44) and the use of first-generation TDI is equivalent [^5] In the SEA, the second-generation TDI signal $\tilde{X}_{2}=\left(1-e^{8\pi ifL}\right)\tilde{X}_{1}$, while the noise PSDs also differ by a multiplicative factor $4\sin^{2}(4\pi fL)=|1-e^{8\pi ifL}|^{2}$. to the use of second-generation TDI. Since the SEA holds to good accuracy for our interested soft memory signals, we consider only the first-generation TDI observables in this paper.
For a single LISA-like detector in the SEA, it is convenient to parameterize the triad $\{\mathbf{a,b,\hat{k}}\}$ in the detector frame defined in the left panel of Fig. 8, where we choose $\mathbf{a}=-\mathbf{e}_{\phi}$, $\mathbf{b}=-\mathbf{e}_{\theta}$ and $\hat{\mathbf{k}}=-\mathbf{e}_{r}$. The positions of the SC ($I\in\{1,2,3\}$) in this detector frame are given by
$\left(\begin{matrix}x_{I}\\ y_{I}\\ z_{I}\end{matrix}\right)=\frac{L}{\sqrt{3}}\left(\begin{matrix}\cos 2\sigma_{I}\\ \sin 2\sigma_{I}\\ 0\end{matrix}\right),\quad\sigma_{I}=\frac{\pi}{4}+I\frac{2\pi}{3},$ (46)
with the guiding center $\mathbf{X=0}$.
For the long-term observation using a single detector or the joint observation using multiple detectors, we parameterize $\{\mathbf{a,b,\hat{k}}\}$ in the solar-system-barycentric (SSB) ecliptic coordinate frame, as depicted in the right panel of Fig. 8. In the equal-armlengths approximation, the positions of SC in the SSB frame can be obtained as
$x_{Ii}^{\prime}=X^{\prime}_{i}(\eta)+R_{ij}(\eta)\,x_{Ij},$ (47)
with the guiding center $\mathbf{X}^{\prime}=R(\cos\eta,\sin\eta,0)$ and $R_{ij}$ a rotation matrix that depends on a phase constant $\Phi_{0}$ (corresponding to $\eta_{0}+\xi_{0}$ in [61]):
$R_{ij}(\eta)=\frac{1}{8}\left(\begin{matrix}\sqrt{3}(s_{2\eta-\Phi_{0}}+3s_{\Phi_{0}}-c_{2\eta-\Phi_{0}}+3c_{\Phi_{0}}&-s_{2\eta-\Phi_{0}}-\sqrt{3}c_{2\eta-\Phi_{0}}-3s_{\Phi_{0}}+3\sqrt{3}c_{\Phi_{0}}&-4\sqrt{3}c_{\eta}\\ -s_{2\eta-\Phi_{0}}-\sqrt{3}c_{2\eta-\Phi_{0}}+3s_{\Phi_{0}}-3\sqrt{3}c_{\Phi_{0}}&\sqrt{3}(3s_{\Phi_{0}}-s_{2\eta-\Phi_{0}})+c_{2\eta-\Phi_{0}}+3c_{\Phi_{0}}&-4\sqrt{3}s_{\eta}\\ 2\sqrt{3}c_{\eta-\Phi_{0}}-6s_{\eta-\Phi_{0}}&2\sqrt{3}s_{\eta-\Phi_{0}}+6c_{\eta-\Phi_{0}}&4\\ \end{matrix}\right),$ (48)
where $c_{z}\equiv\cos z$ and $s_{z}\equiv\sin z$. Note that the coordinates of $\mathbf{e}_{z}$ in the SSB frame is $(-\sqrt{3}\cos\eta,-\sqrt{3}\sin\eta,1)/2$. For the description of GW polarizations, we adopt the standard choice [58]: $\mathbf{a}=-\mathbf{e}_{\phi^{\prime}}$, $\mathbf{b}=-\mathbf{e}_{\theta^{\prime}}$ and $\hat{\mathbf{k}}=-\mathbf{e}_{r^{\prime}}$, with $(\theta^{\prime},\phi^{\prime})=(\pi/2-\beta,\lambda)$, $\beta$ and $\lambda$ being the ecliptic latitude and longtitude of the source direction. The explicit relation between $(\theta,\phi)$ and $(\beta,\lambda)$ is given by $\beta=\arcsin s_{\beta}$, $\lambda=\Theta(c_{\lambda})\,\arcsin s_{\lambda}+\Theta(-c_{\lambda})\,(\pi-\arcsin s_{\lambda})$, with
$\left(\begin{matrix}c_{\beta}c_{\lambda}\\ c_{\beta}s_{\lambda}\\ s_{\beta}\end{matrix}\right)=-\frac{1}{8}\left(\begin{matrix}s_{\theta}\left[-\sqrt{3}(s_{2\eta-\Phi_{0}-\phi}+3s_{\Phi_{0}+\phi})+c_{2\eta-\Phi_{0}-\phi}-3c_{\Phi_{0}+\phi}\right]+4\sqrt{3}c_{\eta}c_{\theta}\\ s_{\theta}\left(s_{2\eta-\Phi_{0}-\phi}+\sqrt{3}c_{2\eta-\Phi_{0}-\phi}-3s_{\Phi_{0}+\phi}+3\sqrt{3}c_{\Phi_{0}+\phi}\right)+4\sqrt{3}s_{\eta}c_{\theta}\\ 2\left[s_{\theta}\left(3s_{\eta-\Phi_{0}-\phi}-\sqrt{3}c_{\eta-\Phi_{0}-\phi}\right)-2c_{\theta}\right]\end{matrix}\right).$ (49)
IV.2 Soft memory signals
Now we turn to the TDI response of a single LISA-like detector to the soft memory events. From Eqs. (10), (12) and (13), the general parametrization of the frequency-domain waveform $\tilde{h}(f)=\tilde{h}_{+}-i\tilde{h}_{\times}$ is
$\tilde{h}=\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi ift_{*}},$ (50)
for the displacement memory,
$\tilde{h}=-\frac{\dot{H}e^{-2i\psi}}{(2\pi f)^{2}}e^{2\pi ift_{*}},$ (51)
for the velocity memory, and
$\tilde{h}=O\,e^{-2i\psi}e^{2\pi ift_{*}},$ (52)
for the integrated-displacement memory, with $H,\dot{H},O>0$ being the memory amplitudes, $\psi\in[0,\pi)$ the polarization angle, and $t_{*}$ the burst arrival time. Note that $\psi$ depends on the choice of polarization tensors (3). Additionally, the sky location of the source is given by $\{\theta,\phi\}$ in the detector frame, and $\{\beta,\lambda\}$ in the SSB frame. The former will be used in the following discussion of single-detector response.
Figure 9: Comparison between analytical and numerical spectrum of $A_{1}(t)$ for the arctan model (7) at LISA, with parameters $H=10^{-19}$, $T=0.05\,\text{yr}$, $t_{*}=0.5T$, $\{\beta,\lambda,\psi\}=\{-1.4,2.3,\pi/3\}$, and $f_{*}\in\{10^{-5},0.01,1,5\}\,\text{Hz}$. The analytical spectrum (black dashed line) is obtained in the SEA, with the soft limit indicated by the blue dashed line. The numeric spectrum (colored solid line) is computed using fastlisaresponse with the equal-armlength orbit and a sampling frequency of $1\,\text{Hz}$. The vertical dashed lines correspond to integer multiples of $f=1/(2L)$. The time-domain waveforms are shown in the figure inset.
In the SEA, the frequency-domain signals $\{\tilde{A}_{1},\tilde{E}_{1},\tilde{T}_{1}\}$ can be calculated using Eqs. (42) and (45). The signals are degenerate under the transformation from $\{\theta,\psi\}$ to $\{\theta^{\prime},\psi^{\prime}\}=\{\pi-\theta,\pi-\psi\}$, and merely changes sign under the transformation from $\psi$ to $\psi^{\prime}=\psi\pm\pi/2$. In the low-frequency approximation (44), there is an additional degeneracy under $\phi\leftrightarrow\phi+\pi$. Moreover, the signals in the SEA at $f\to 0$ depend only on four parameters, while a soft memory signal has five parameters in total. The measurement of a soft memory event with a single detector is therefore possible precisely because the response function deviates from its low-frequency approximation. When this deviation is small, the signal suffers from a stronger degeneracy.
Figure 10: Normalized logarithmic spectral density of $\rho_{A+E}^{2}$ (top panel) and the corresponding signal spectrum (bottom panel) at Taiji and BBO, for a soft displacement-memory signal with $\{\phi,\theta,\psi\}=\{0,0,0\}$.
Figure 11: Normalized logarithmic spectral density of $\rho_{A+E}^{2}$ (top panel) and the corresponding signal spectrum (bottom panel) at Taiji and BBO, for a soft velocity-memory signal with $\{\phi,\theta,\psi\}=\{0,0,0\}$.
Figure 12: Normalized logarithmic spectral density of $\rho_{A+E}^{2}$ at Taiji and BBO, for a soft integrated-displacement-memory signal with $\{\phi,\theta,\psi\}=\{0,0,0\}$.
For the LISA detector, we verified that the analytical approximation matches the time-domain computation of fastlisaresponse [58], up to an overall sign. The static-detector approximation in fact remains good when the deviation from the soft limit is large. This is shown in Fig. 9 for the arctan model (7) with different transition time scale $1/f_{*}$. It is seen that for $f_{*}=1\,\text{Hz}$, the spectrum is already close to the soft waveform at $f<0.01\,\text{Hz}$. For $f_{*}=10^{-5}\,\text{Hz}$, the soft limit does not provide a good approximation at $f>10^{-6}\,\text{Hz}$, but the SEA remains accurate. The situation is expected to be the same for a generic waveform with displacement memory.
IV.3 Optimal signal-to-noise ratio
We proceed to estimate the detectability of soft memory events through matched filtering. Since the $\{A,E,T\}$ channels are orthogonal in the SEA, the noise-weighted inner product between two time-domain signals $U_{a}(t)$ and $U_{b}(t)$ is
$\langle a|b\rangle\equiv\sum_{U=A,E,T}\int_{f_{\text{min}}}^{f_{\text{max}}}df\frac{4\,\text{Re}\left[\tilde{U}_{a}(f)\,\tilde{U}_{b}^{*}(f)\right]}{S_{U}(f)},$ (53)
where $S_{U}(f)$ denotes the one-sided power spectral density (PSD) of the noise in the $U$-th channel, which is assumed to be wide-sense stationary. In practice, the continuous integral is replaced by a discrete sum with frequency step $\Delta f=f_{\text{min}}$, but their difference is small if $f_{\text{min}}$ is sufficiently low; in this section, we use the continuous integral. The noise PSD takes the form [70, 4, 45]:
[…]
$+8\left[2+\cos(2\pi fL)\right]\,S_{\text{oms}}.$
For LISA, Taiji and TianQin, we use the optical metrology noise $S_{\mathrm{oms}}(f)$ and test mass acceleration noise $S_{\text{acc}}(f)$ given by
$S_{\mathrm{oms}}$ $=\left(\frac{2\pi fs_{\text{acc}}}{c}\right)^{2}\left[1+\left(\frac{2\,\mathrm{mHz}}{f}\right)^{4}\right]\text{s},$ (55)
$S_{\mathrm{acc}}$ $=\left(\frac{s_{\text{acc}}}{2\pi fc}\right)^{2}\left[1+\left(\frac{4\,\mathrm{mHz}}{10f}\right)^{2}\right]\left[1+\left(\frac{f}{8\,\mathrm{mHz}}\right)^{4}\right]\text{s}.$ (56)
The values of $s_{\text{oms}}$ and $s_{\text{acc}}$ are listed in Table 1. For the BBO [21, 20], we use
$S_{\mathrm{oms}}=\frac{2\times 10^{-34}\,\text{s}}{(3L/\text{m})^{2}},\quad S_{\mathrm{acc}}=\frac{9\times 10^{-34}\,\text{s}}{(2\pi f/\text{Hz})^{4}(3L/\text{m})^{2}}.$ (57)
In its proposed design, the first stage of BBO comprises a single LISA-like constellation, and the second stage consists of 12 SC in three constellations. To be conservative, here we consider the first stage of BBO. We also neglect the impact of Galactic confusion noise and other signal components in the data streams, which can only be addressed through global fitting.
For a given signal $U_{a}(t)$, the optimal signal-to-noise ratio (SNR) $\rho$ of a matched filtering is given by
$\rho^{2}=\langle a|a\rangle=\sum_{U}\int d\ln f\left[\frac{4f|\tilde{U}_{a}|^{2}}{S_{U}(f)}\right].$ (58)
In this paper, we do not include the T channel, which is much weaker than the A and E channels and is significantly affected by unequal armlengths at low frequencies; see Appendix A for further discussion. The resulting SNR is denoted by $\rho_{A+E}$, which turns out to be nearly independent of $\phi$. Note that $S_{A,E}=0$ at $f=1/(2L)$, but no divergence occurs because the factor $\left|1-e^{4\pi ifL}\right|^{2}=4\sin^{2}(2\pi fL)$ is canceled by the corresponding factor in the signal in Eq. (43).
The logarithmic spectral density of soft displacement-memory and velocity-memory signal for $\phi=\theta=\psi=0$ are shown in Figs. 10 and 11, respectively. The spectral density is normalized such that the peak value equals unity. We also show results obtained using the low-frequency approximation of the response function as dashed lines; these correspond to the flat signal spectrum at the low-frequency end in the bottom panel. As can be seen, a significant portion of the spectrum cannot be described using the low-frequency approximation in the case of displacement memory. Meanwhile, the low-frequency approximation is much better for Taiji in the case of velocity memory, indicating a strong degeneracy of the signal.
To quantify the spectral range where the SNR receives the most contribution, we introduce the frequencies $f_{\text{I}}$, $f_{\text{II}}$ and $f_{0}$ as follows. For $\theta=\psi=\phi=0$, $\rho^{2}_{A+E}(f_{\text{min}}=10^{-5}\,\text{Hz},f_{\text{max}}=2f_{\text{I}})/\rho^{2}_{A+E}(f_{\text{min}}=10^{-5}\,\text{Hz},f_{\text{max}}=f_{\text{I}})-1\approx 10^{-2}$, $\rho^{2}_{A+E}(f_{\text{min}}=10^{-5}\,\text{Hz},f_{\text{max}}=f_{\text{II}})\approx\frac{1}{2}\rho^{2}_{A+E}(f_{\text{min}}=10^{-5}\,\text{Hz},f_{\text{max}}=f_{\text{I}})$, and $\rho^{2}_{A+E}(f_{\text{min}}=f_{0}/2,f_{\text{max}}=f_{\text{I}})/\rho^{2}_{A+E}(f_{\text{min}}=f_{0},f_{\text{max}}=f_{\text{I}})-1\approx 10^{-2}$. Thus, the maximum SNR for $\phi=\theta=\psi=0$ is largely determined by the spectrum within $f_{\text{0}},f_{\text{I}}]$, and half of the maximum squared SNR is contributed by the frequency range $f_{\text{0}},f_{\text{II}}]$. The SNR is nearly unaffected by the signal duration once it exceeds $1/f_{0}$. We list the values of $f_{\text{0,II,I}}$ for LISA, Taiji, TianQin, and BBO in Table 2. As can be seen, the low-frequency part of the spectrum is more important for the velocity memory. The situation is reversed in the case of integrated-displacement memory, as shown in Fig. 12. In the following, we set $f_{\text{min}}=f_{0}$ and $f_{\text{max}}=f_{\text{I}}$, unless stated otherwise.
The angular dependence of the SNR exhibits the following symmetry: $\rho_{A+E}(\phi,\theta,\psi)=\rho_{A+E}(\phi,\pi-\theta,\pi-\psi)=\rho_{A+E}(\phi,\theta,\psi+\pi/2)=\rho_{A+E}(\phi+\pi,\theta,\psi)$. To a good approximation, we also have $\rho_{A+E}(\phi,\theta,\psi)\approx\rho_{A+E}(\theta,\psi)\approx\rho_{A+E}(\pi-\theta,\psi)$. So the angular dependence of the SNR is almost fixed by $\rho^{2}_{A+E}(\theta,\psi)$ within $\psi\in[0,\pi/2]$ and $\theta\in[0,\pi/2]$. The result is roughly the same for all considered detectors and is shown for Taiji in Fig. 13. The results for displacement and velocity memory are also found to be similar. The SNR is maximized at $\theta\in\{0,\pi\}$, where it is independent of $\psi$. The SNR is minimized at $\theta=\pi/2$, $\psi\in\{\pi/4,3\pi/4\}$, where it is approximately $1.6\%$ of the maximum SNR. For a given $\theta$, the SNR is maximized at $\psi\in\{0,\pi/2,\pi\}$, and minimized at $\psi\in\{\pi/4,3\pi/4\}$; while for a given $\psi$, the SNR is maximized at $\theta\in\{0,\pi\}$, and minimized at $\theta=\pi/2$. In Table 2, we list the minimum memory amplitudes $H_{10}^{\text{min}}$ and $\dot{H}_{10}^{\text{min}}$ for $\rho_{A+E}=10$ at the considered detectors, which correspond to the events with $\theta=\psi=0$.
It is also useful to examine the SNR averaged over $\{\phi,\theta,\psi\}$, which provides a measure of the “mean” signal strength. Generically, a linearly polarized GW can be written as $h_{ij}(t)=\left(c_{2\psi}\,e_{ij}^{+}+s_{2\psi}\,e_{ij}^{\times}\right)\mathcal{H}(t)$, and the signal in the static-detector approximation takes the form: $\tilde{U}(f)=\mathcal{F}^{U}_{ij}(f)\,\tilde{h}_{ij}(f)=\tilde{U}(\psi=0)\,c_{2\psi}+\tilde{U}(\psi=\pi/4)\,s_{2\psi}$. The averaged version of Eq. (58) is then given by
$\left\langle\rho^{2}\right\rangle$ $\equiv\int\frac{\sin\theta\,d\theta\,d\phi\,d\psi}{8\pi^{2}}\langle a|a\rangle$ (59)
$=\int\frac{\sin\theta\,d\theta\,d\phi}{4\pi}\frac{1}{2}\left(\langle a|a\rangle_{\psi=0}+\langle a|a\rangle_{\psi=\pi/4}\right)$
$=\sum_{U}\sum_{\lambda=+,\times}\int df\,\mathcal{F}^{U}_{\lambda}\,\left|\tilde{h}_{\lambda}(\psi=0)\right|^{2},$
where
$\mathcal{F}^{U}_{\lambda}(f)\equiv\int\frac{d\Omega_{2}}{8\pi}\,\left|\mathcal{F}^{U}_{ij}e_{ij}^{\lambda}\right|^{2}.$ (60)
Note that in the SEA, the low-frequency limit is [45]
$\lim_{f\to 0}\sum_{\lambda}\mathcal{F}_{\lambda}^{A,E}\to\frac{576}{5}(fL\pi)^{4}.$ (61)
In Table 2, we list the averaged memory amplitudes $\langle H\rangle_{10}$ and $\langle\dot{H}\rangle_{10}$ for $\sqrt{\langle\rho_{A+E}^{2}\rangle}=10$ at the considered detectors. We find that the values of $H_{10}$ and $\dot{H}_{10}$ of the same order of magnitude for LISA, Taiji and TianQin, whereas BBO can achieve significantly higher sensitivity to the soft displacement-memory signal, provided it occurs at $f\lesssim 1\,\text{Hz}$.
Detector $s_{\text{oms}}$ [fm] $s_{\text{acc}}$ [pm] $L$ [Gm] $1/(2L)$ [Hz]
LISA [3, 4] 3 15 2.5 0.06
Taiji [50] 3 8 3 0.05
TianQin [65] 1 1 0.17 0.88
BBO [20] $\times$ $\times$ 0.05 3
Table 1: Parameters used for the considered detectors.
Figure 13: Angular distribution of $\rho_{A+E}(\theta,\psi)/\rho_{A+E}(0,0)$ at Taiji for $\phi=0$ and given $H$ ($\dot{H}$) in the case of displacement (velocity) memory, shown as the solid (dashed) contours.
 LISA  Taiji  TianQin  BBO
$H^{\text{min}}_{10}/10^{-20}$ 1.3 0.69 2.0 $2.6\times 10^{-4}$
$\langle H\rangle_{10}/10^{-20}$ 2.0 1.1 3.3 $3.9\times 10^{-4}$
$f_{0}$ [mHz] 0.56 0.71 1.1 127
[…]
$f_{\text{II}}$ [mHz] 5.1 6.4 11 1220
$f_{\text{I}}$ [Hz] 0.032 0.035 0.25 25
$\dot{H}^{\text{min}}_{10}/10^{-22}$ [Hz] 2.7 1.8 8.5 0.13
$\langle\dot{H}\rangle_{10}/10^{-22}$ [Hz] 4.2 2.9 13 0.19
$f_{0}$ [mHz] 0.13 0.16 0.19 51
[…]
$f_{\text{II}}$ [mHz] 2.1 2.7 3.9 462
$f_{\text{I}}$ [Hz] 0.010 0.013 0.019 2.0
Table 2: $f_{\text{0,I,II}}$, and the minimum and averaged memory amplitudes for $\rho_{A+E}=10$ at the considered detectors, in the cases of displacement (top rows) and velocity memory (bottom rows). The minimum memory amplitude threshold and $f_{\text{0,I,II}}$ are obtained at $\phi=\theta=\psi=0$, whereas the averaged memory amplitude threshold is obtained using the averaged response function.
IV.4 Distinguishability of real linear and quadratic corrections to the soft displacement-memory signal
As discussed in Sec. II.4, there is necessarily a correction to the soft waveform of displacement memory in the form of Eq. (15). However, this cannot be faithfully distinguished unless the SNR is sufficiently large. As concrete examples, in this section we estimate the distinguishability of real linear and quadratic corrections, corresponding to $C(f)=a_{0}f$ and $C(f)=a_{1}f^{2}$, respectively, with $a_{0,1}\in\mathbb{R}$. In the following, we consider the cases of linear and quadratic corrections separately.
The difference between two waveforms $h_{1}$ and $h_{2}$ is reflected in their mismatch, defined as
$\mathcal{M}=1-\frac{\langle h_{1}|h_{2}\rangle}{\sqrt{\langle h_{1}|h_{1}\rangle\langle h_{2}|h_{2}\rangle}}.$ (62)
A rough estimate for the condition of distinguishability is [13]
$\rho>\rho_{\text{crit}}=\sqrt{\frac{D}{2\mathcal{M}}},$ (63)
where $D$ is the number of parameters in the waveform template. In the present case, we choose $h_{1}$ ($h_{2}$) to be the waveform with (without) the correction, and take $D=6$. Note that for a generic correction factor $C(f)\in\mathbb{C}$, $\mathcal{M}$ depends only on the integrals $\sum_{U}\int d\ln f\left[\frac{4f|\tilde{U}_{a}|^{2}}{S_{U}(f)}\right]\text{Re}\,C$ and $\sum_{U}\int d\ln f\left[\frac{4f|\tilde{U}_{a}|^{2}}{S_{U}(f)}\right]|C|^{2}$. Fig. 14 shows the estimated critical values of $a_{0,1}$ distinguishable by Taiji and BBO as functions of the SNR, which are insensitive to $\{\phi,\theta,\psi\}$. In the next section, we will examine the impact of such corrections to the parameter estimation using the uncorrected soft-waveform template in concrete examples.
Figure 14: Estimated maximum and minimum values of $a_{0}$ ($a_{1}$) distinguishable by Taiji and BBO, shown as functions of $\rho_{A+E}$, for linear (quadratic) corrections. The solid lines correspond to $\theta=\psi=0$, while the black dashed line shows the result obtained using the averaged response function of Taiji. For BBO, the values of $10^{2}a_{0}$ and $10^{4}a_{1}$ are shown.
V Parameter estimation simulation
Once a soft memory signal is identified through matched filtering, further measurements through Bayesian parameter estimation can be performed. Given the data $d$ and a model $\Lambda$ with parameters $\boldsymbol{\xi}$, the posterior probability distribution, according to Bayes’ rule, is
$P(\boldsymbol{\xi}|d,\Lambda)=\frac{p(\boldsymbol{\xi}|\Lambda)\,\mathcal{L}}{\mathcal{Z}},$ (64)
where $\mathcal{L}=p(d|\boldsymbol{\xi},\Lambda)$ is the likelihood, $\mathcal{Z}=p(d|\Lambda)=\int d\boldsymbol{\xi}\,p(\boldsymbol{\xi}|\Lambda)\,\mathcal{L}$ is the evidence, and $p(\boldsymbol{\xi}|\Lambda)$ is the prior. The likelihood is taken to be
$\mathcal{L}\propto\exp\left(-\frac{1}{2}\langle d-h\mid d-h\rangle\right).$ (65)
In the present case, the data containing the signal $s(t)$ in a TDI channel is given by $d(t)=s(t)+n(t)$, with $n(t)$ being the noise. Idealistically, the noise is wide-sense stationary, described by the PSD (54).
In the limit of high SNR, the measurement precision can be estimated using the Fisher information matrix (FIM). For a model $h_{\boldsymbol{\xi}}(t)$ with parameters $\boldsymbol{\xi}=\{\xi_{i}\}_{i}$, the FIM evaluated at the true parameter values $\boldsymbol{\xi}=\boldsymbol{\hat{\xi}}$ is $\Gamma_{ij}\equiv\left\langle{\partial h}/{\partial\xi_{i}}\mid{\partial h}/{\partial\xi_{j}}\right\rangle_{\boldsymbol{\xi}=\boldsymbol{\hat{\xi}}}$, and the inferred covariance matrix of $\boldsymbol{\xi}$ is estimated to be $\Sigma_{ij}=(\Gamma^{-1})_{ij}$, with the root-mean-squared error of $\xi_{i}$ given by $\sigma_{i}=\sqrt{\Sigma_{ii}}$. Since $t_{*}\in(0,T)$, we use $\tau_{*}=t_{*}/T\in(0,1)$ as a parameter. We then consider a five-parameter model with $\boldsymbol{\xi}=\{H,\tau_{*},\phi,\cos\theta,\psi\}$. The angular resolution of the source sky location, given by the solid-angle error corresponding to the $1\sigma$ error ellipse, can be estimated as
$\sigma_{\Omega}=2\pi\sqrt{\Sigma_{c_{\theta}c_{\theta}}\Sigma_{\phi\phi}-\Sigma_{c_{\theta}\phi}^{2}}=2\pi\sqrt{\Sigma_{s_{\beta}s_{\beta}}\Sigma_{\lambda\lambda}-\Sigma_{s_{\beta}\lambda}^{2}}~,$
which is invariant under coordinate rotation. The result for $\phi=0$ is shown in Fig. 15. It should be noted that the result is $\phi$-dependent, and the estimated errors are proportional to $1/H$, while $\sigma_{\tau_{*}}$ is also proportional to $1/T$. Due to the symmetry of the signal, we show only $\psi\in[0,\pi/2]$ and $\theta\in[0,\pi/2]$. Notably, the FIM estimate of the measurement error increases drastically as $\theta\to\pi/2$. The reason is that when $\theta=\pi/2$, only $h_{+}$ contributes to the signal, hence $H$ and $\psi$ are degenerate; additionally, $\phi$ and $\psi$ are degenerate at $\theta\in\{0,\pi\}$. For a given $\theta$, the estimated error is minimized at $\psi\in\{0,\pi/2\}$. We also observe that the FIM estimate of the measurement error is larger for velocity memory at roughly the same SNR, possibly due to its stronger degeneracy discussed in Sec. IV.2.
Figure 15: FIM estimate of $\sigma_{\Omega}/(4\pi)$ for a soft memory signal at Taiji, in the cases of displacement (top panel) and velocity (bottom panel) memory. The memory amplitudes are set to $H=10^{-19}$ and $\dot{H}=2.6\times 10^{-21}\,\text{Hz}$, so that $\rho_{A+E}$ at $\theta=\psi=0$ is identical. The other parameters used are $\phi=0$, $\tau_{*}=0.3$ and $T=10^{-3}\,\text{yr}$.
However, the FIM estimate cannot be reliably trusted at low SNR, which is likely a realistic regime. In the following, we present results from Bayesian parameter-estimation simulations, focusing on the case of displacement memory. For sampling the posterior, we use the MCMC sampler emcee [34] and plot the posterior distributions using corner [35]. For concreteness, we consider the measurements using Taiji, LISA-Taiji network (described in Fig. 8) and BBO. For LISA-Taiji, we set $\eta_{\text{Taiji}}=\eta_{\text{LISA}}+2\pi/9$, $\Phi_{0}=0$, and $\eta_{\text{LISA}}=\pi/4$; the log-likelihood is taken as the sum of those of LISA and Taiji. We use the parameter set $\boldsymbol{\xi}=\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$ for LISA-Taiji, and $\boldsymbol{\xi}=\{H,\tau_{*},\phi,\cos\theta,\psi\}$ for a single BBO detector. The priors are taken to be uniform, as listed in Table 3.
Parameter Lower Bound Upper Bound
$\log_{10}H$ $\log_{10}\hat{H}-2$ $\log_{10}\hat{H}+2$
$\log_{10}(\dot{H}/\text{Hz})$ $\log_{10}(\hat{\dot{H}}/\text{Hz})-2$ $\log_{10}(\hat{\dot{H}}/\text{Hz})+2$
$\tau_{*}$ 0 1
$\psi$ 0 $\pi$
$\phi$ 0 $2\pi$
$\cos\theta$ 1-1 1
$\Psi$ 0 $\pi$
$\lambda$ 0 $2\pi$
$\sin\beta$ 1-1 1
Table 3: Lower and upper bounds for the parameters in the uniform distribution. The parameters $\{\psi,\phi,\cos\theta\}$ are defined in the detector frame for a single detector, while the parameters $\{\Psi,\lambda,\sin\beta\}$ are defined in the SSB frame (see Sec. IV.1). $\psi$ and $\Psi$ are the polarization angles in the detector and SSB frames, respectively.
[…]
Figure 21: Spectra of $2\pi f|\tilde{h}(f)|/H=1+C(f)$ in various scenarios. The solid and dashed lines represent the real linear and quadratic corrections considered in Fig. 19, respectively. The blue lines correspond to $|a_{0}|=50\,\text{Hz}^{-1}$ or $|a_{1}|=20000\,\text{Hz}^{-2}$, the red lines correspond to $|a_{0}|=10\,\text{Hz}^{-1}$ or $|a_{1}|=10000\,\text{Hz}^{-2}$. The gray and black dot-dashed lines correspond to the waveforms of compact binary scattering and merger events considered in Fig. 20, respectively.
Fig. 16 shows a simulated measurement result by LISA-Taiji in the high-SNR regime, with $\hat{H}=10^{-19}$. As can be seen, $\tau_{*}$ is measured with significantly higher precision (this is partly because the considered signal duration is relatively long), while the memory amplitude $H$ is better constrained than the angular parameters $\{\lambda,\beta,\Psi\}$. Joint observations with LISA and Taiji is found to greatly improve the measurement accuracy compared with observations using Taiji alone, even though the SNR is similar. We also test the stochasticity of the noise injection and find that, while it does not significantly affect the results, it introduces an overall bias. In the absence of noise injection, the results agree well with the FIM estimation.
The lower-SNR regime is examined in Fig. 17. We find that even for $\rho\sim 10$, observations with Taiji alone can still effectively constrain the parameters, in particular $H$ and $\tau_{*}$. Notably, the FIM substantially overestimates the measurement error when the SNR is low. If the sky location of the source is known a priori (from other observations), the measurement precision improves dramatically. For the soft velocity-memory signal with the same SNR, the measurement errors are slightly larger. The impact of source direction on the measurement precision is examined in the upper panel of Fig. 17, where the source colatitude $\theta$ in the detector frame of Taiji is varied while the other parameters and the SNRs are kept fixed at 13.7. The measurement errors are found to be of the same order of magnitude, although $\sigma_{s_{\beta}}$ is relatively large for $\hat{\theta}=0.01\pi$, likely due to the signal degeneracy at $\theta=0$.
The bottom panel of Fig. 18 shows the simulated measurement results by a single BBO detector at different SNRs. We again find that the measurement precision of $H$ and $\tau_{*}$ is better when the SNR is sufficiently high. Compared with measurements using Taiji, the results agree better with the FIM estimates for $\rho\sim 10$, possibly due to weaker parameter degeneracies of the signal at BBO.
Figs. 19-20 present example results of parameter estimation for corrected soft waveforms using an uncorrected template. In Fig. 19, we show the relative measurement errors of a soft displacement-memory signal with real linear and quadratic corrections (as discussed in Sec. IV.4), i.e., $\tilde{h}=\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi ift_{*}}[1+C(f)]$ with $C(f)=a_{0}f+a_{1}f^{2}$. To isolate the intrinsic parameter-estimation performance, we omit noise injection. The considered correction factors are relatively large and are plotted in Fig. 21. We find that these corrections only have a minor impact on the measurement of $\tau_{*}$, but they introduce clear biases in $\log_{10}H$. When the correction is moderate, however, the bias remains modest, staying within the $2\sigma$ deviation. For the other parameters, which have considerably lower measurement precision and show visible bias even in the absence of corrections, the results are not significantly affected, although the biases tend to increase.
As realistic examples, we simulate the measurement of a stellar-mass compact binary scattering event with the Taiji detector and a quasi-circular stellar-mass BBH merger event with the BBO detector. For the former, we use the waveform in Fig. 4, and for the latter, the $q=1$ non-spinning waveform in Fig. 7 (taking into account the comic expansion). As discussed previously, both waveforms can be approximated by soft waveforms of displacement memory with real correction factors $C(f)$, shown in Fig. 21. As above, a soft-waveform template is employed to perform the parameter estimation. The posterior distributions of the memory amplitude and polarization angle, together with the detailed parameter settings, are shown in Fig. 20. In both examples, we find that the memory parameters can be effectively constrained, with the amplitude $H$ measured with a higher precision.
VI Stochastic soft memory signals
A data stream of sufficiently long duration may contain multiple, or even a large number of, events with memory. In the latter case, the detector effectively observes a stochastic GW background (SGWB) of soft memory signals [99]. Even if the individual events cannot be measured, the SGWB may still be detectable if the event rate is sufficiently large. This could be the case, e.g., for the SGWB from compact binary scattering [36, 59]. The SGWB from compact binary mergers will also include a contribution from null memory, although it is expected to be sub-leading compared with the contribution from oscillatory GWs emitted by the low-frequency sources within the same population. In this section, we examine the signal PSD of an idealized SGWB of soft displacement-memory signals at the LISA-like detectors and assess its detectability. We will also briefly comment on the SGWB of soft velocity-memory signals.
Generically, the GW can be written as
$h_{ij}(t,\mathbf{x})=\sum_{\lambda}\int_{-\infty}^{\infty}df\,d\Omega_{2}\,\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,e_{ij}^{\lambda}(\hat{\mathbf{k}})\,e^{-2\pi if(t-\hat{\mathbf{k}}\cdot\mathbf{x})}.$ (66)
If the SGWB is stationary, isotropic and unpolarized (such an SGWB will henceforth be referred to as “idealized”), the two-point statistics satisfy [73]
$\left\langle\tilde{h}_{\lambda}(f,\hat{\mathbf{k}})\,\tilde{h}^{*}_{\lambda^{\prime}}(f^{\prime},\hat{\mathbf{k}}^{\prime})\right\rangle=\frac{S_{h}(f)}{16\pi}\,\delta(f-f^{\prime})\,\delta^{2}(\hat{\mathbf{k}},\hat{\mathbf{k}}^{\prime})\,\delta_{\lambda,\lambda^{\prime}},$ (67)
[…]
$\Omega_{\text{gw}}(f)=\frac{1}{\rho_{\text{crit}}}\frac{d\rho_{\text{gw}}}{d\ln f}=\frac{2\pi^{2}}{3H_{0}^{2}}f^{3}S_{h}(f),$ (68)
with $H_{0}$ and $\rho_{\text{crit}}=3H_{0}^{2}/(8\pi)$ being the Hubble parameter and critical density today.
[…]
$=\frac{S_{h}}{8\pi}\int d\Omega_{2}\,\Lambda_{ij,kl}$
$=\frac{S_{h}}{10}\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ij}\delta_{kl}\right),$
where we used the TT polarization sum $\sum_{\lambda}e_{ij}^{\lambda}\,e_{kl}^{\lambda*}=2\Lambda_{ij,kl}$, with $\Lambda_{ij,kl}$ defined in Eq. (17). Thus $\mathcal{P}_{ij}=\frac{1}{5}\left(1+\frac{1}{3}\delta_{ij}\right)S_{h}$, $\mathcal{P}_{i,j=i}=\frac{4}{15}S_{h}$, and
$S_{h}=\frac{1}{2}\sum_{i,j}\mathcal{P}_{ij}.$ (70)
For a LISA-like detector, the $U$-channel TDI signal in the static-detector approximation is given by $\tilde{U}(f)=\mathcal{F}^{U}_{ij}(f)\,\tilde{h}_{ij}(f)$. Using Eq. (69), the one-sided PSD of $U(t)$ is thus
$P_{U}(f)=\frac{S_{h}}{4\pi}\int d\Omega_{2}\,\Lambda_{ij,kl}\,\mathcal{F}^{U}_{ij}\,\mathcal{F}^{U*}_{kl}=S_{h}\sum_{\lambda}\mathcal{F}^{U}_{\lambda},$ (71)
where $\mathcal{F}^{U}_{\lambda}(f)$ is the averaged response function defined in Eq. (60).
Figure 22: Comparison between the simulated 1-year SGWB of soft displacement-memory signals at LISA with analytical prediction. The top panel shows the time-domain waveform $h_{33}(t)$; the middle panel shows the SGWB spectrum; the bottom panel shows the PSD and CSD spectra of TDI signals.
 LISA  Taiji  TianQin  BBO
$\Delta_{10}/10^{-47}$ 5.6 1.8 2.2 $5.0\times 10^{-6}$
Table 4: Value of the displacement memory SGWB parameter $\Delta$ for $\rho_{A+E}=10$ and $T=1\,\text{yr}$ at the considered detectors.
Figure 23: Normalized logarithmic spectral density of $\rho_{A+E}^{2}$ at LISA, Taiji, TianQin and BBO, for an idealidealized SGWB of soft displacement-memory signals.
Figure 24: Simulated 1-year SGWB of soft velocity-memory signals.
For an ensemble of soft displacement-memory signals, the resulting $h_{ij}(t)$ is a superposition of step functions with different heights $\mathcal{H}$ and step positions $t_{*}$. Such a time-domain waveform is similar to a Brownian motion and is not wide-sense stationary. Nevertheless, the Welch PSD estimator turns out to be
$\mathcal{P}_{ij}\approx\frac{2}{T}\frac{\sum\mathcal{H}^{2}}{(2\pi f)^{2}}=\frac{N}{T}\frac{2\langle\mathcal{H}^{2}\rangle}{(2\pi f)^{2}},$ (72)
(which is compatible with [99]) where $N$ is the total number of steps, and we denote the mean and standard deviation of $X$ by $\langle X\rangle$ and $\sigma_{X}$, respectively.
Now we consider an ensemble of mutually-independent soft displacement-memory signals with a uniform distribution of $\{\tau_{*},\lambda,\sin\beta,\Psi\}$ and a random distribution of $H$. In the present case, we refer to such a SGWB as “idealized”. A single event manifests in $h_{33}(t)$ as a step with height
$\mathcal{H}=\kappa H,\quad\kappa=-c^{2}_{\beta}\,c_{2\Psi}.$ (73)
It follows that $\langle\mathcal{H}\rangle=0$, and $\langle\mathcal{H}^{2}\rangle=\sigma_{\kappa}^{2}\left(\langle H\rangle^{2}+\sigma_{H}^{2}\right)$, with $\sigma_{\kappa}^{2}=4/15$. Eqs. (70) and (72) then yield
$S_{h}=\frac{\Delta}{f^{2}},$ (74)
where
$\Delta=\frac{2N}{T}\frac{\langle H\rangle^{2}+\sigma_{H}^{2}}{(2\pi f)^{2}}.$ (75)
The corresponding energy density $d\rho_{\text{gw}}/df$ [defined by Eq. (68)] is equivalent to a simple sum of the spectral energy flux density $\frac{1}{T}\frac{dE_{\text{gw}}}{R^{2}d\Omega_{2}df}=\frac{H^{2}}{8\pi T}$ associated with each memory jump. For example, if $H$ is distributed uniformly in $0,H_{\text{max}}]$, then $\langle H\rangle^{2}+\sigma_{H}^{2}=\frac{1}{3}H_{\text{max}}^{2}$. To verify this description, a time-domain simulation of the SGWB is performed and the first-generation TDI signal at LISA detector is computed using fastlisaresponse with the realistic ESA orbit. The SGWB is constructed from 4000 sources drawn from a uniform distribution of $\{H,\tau_{*},\lambda,\sin\beta,\Psi\}$. We find that the result agrees well with the analytical prediction to $S_{h}$ and the signal PSD given by Eq. (71), as shown in Fig. 22.
In principle, the signal PSD can be measured [45, 93], and the corresponding SNR squared is given by
$\rho^{2}=T\sum_{U}\int_{f_{\text{min}}}^{f_{\text{max}}}df\,\left(\frac{P_{U}}{S_{U}}\right)^{2}.$ (76)
The SNR of the T channel is again found to be small and will be neglected. In Table 4 we list the values of the parameter $\Delta$ in Eq. (74) for $\rho_{A+E}=10$ and $T=1\,\text{yr}$ at the considered detectors. The corresponding logarithmic spectral density of $\rho_{A+E}^{2}$ is shown in Fig. 23. For LISA, Taiji and TianQin, the SNR is roughly contributed by $f\in[10^{-3},0.1]\,\text{Hz}$, while for BBO it is roughly contributed by $f\in[0.1,10]\,\text{Hz}$.
In contrast to the case of displacement memory, the SGWB of soft velocity-memory signals cannot be characterized by a PSD. Fig. 24 shows an example of such a SGWB, constructed from $10^{4}$ sources modeled as an ensemble of ReLU waveforms with $\Delta\dot{h}_{-}=0$ (see Sec. II.2), adopting a uniform distribution of $\{\dot{H},\tau_{*},\lambda,\sin\beta,\Psi\}$. The time-domain waveform resembles a low-frequency deterministic signal. Detecting such a signal component is expected to be particularly challenging.
VII Summary and discussion
As predicted by general relativity, sources of high-frequency GWs can also produce soft signals in detectors sensitive only to low-frequency bands. Such signals take clean, universal forms, with the memory amplitudes as the only intrinsic parameters. We have investigated the prospects for detecting such signals with future space-based interferometers, through observations of individual events and the stochastic background. For a single event, the SNR in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$. Such a signal could potentially be produced by compact binary scatterings, supernovae, Gamma-ray bursts, or certain dark transients. The SNR of a single event in BBO can reach 10 for a displacement memory of amplitude $\sim 10^{-24}$. Such signals are generically expected from stellar-mass compact binary mergers, due to the nonlinear nature of gravity. We examine the parameter-estimation precision for soft displacement-memory signals using simulated Bayesian analyses. To characterize the deviation of the GW spectrum from the zero-frequency limit or an exact soft waveform (e.g., in compact binary mergers), an effective correction factor can be introduced. We examine such corrections in realistic examples of compact binary scattering and quasi-circular BBH mergers, and find that the soft waveform can still serve as an effective template in these cases. A large number of unresolvable soft signals can also form a stochastic gravitational-wave background. For displacement memory, the resulting background is expected to follow a strain power spectral density $S_{h}\propto f^{-2}$. We evaluate the corresponding SNR of the signal power spectrum at the planned detectors considered in this paper. Our results indicate that soft signals from bursts with memory can be promising targets for space-based gravitational-wave detection.
In this paper, we only consider the soft memory signals in general relativity, but they may also arise in additional GW polarization modes [82, 49, 46]. Similar signals could arise from massless or ultralight dark radiation, as discussed in Appendix B. The relevant templates can be readily constructed, and the distinguishability analysis can be performed. Finally, soft memory signals that extend to sufficiently high frequencies can, in principle, be jointly observed by ground-based and space-based detectors, providing complementary information about the spectrum. The possible soft memory signals associated with a GW event observed by a ground-based detector can also be searched for with space-based detectors.
Acknowledgments
Y. L. M. is supported in part by the National Key R & D Program of China under Grant No. 2021YFC2202900, the National Science Foundation of China (NSFC) under Grant No. 12547104 and Gusu Talent Innovation Program under Grant No. ZXL2024363. Y. T is partly supported by NSFC under Grant No. 12547104.
Appendix A Soft memory signals in the T channel
In the main text, the signal in the T channel is neglected. In contrast to $A$ and $E$, the low-frequency spectrum of $T$ is sensitive to the armlength differences and deviates drastically from the result in SEA. We therefore need to consider second-generation TDI variables, e.g.,
$X_{2}=$ $X_{1}+[(D_{13121}y_{12}+D_{131212}y_{21}$ (77)
$+D_{1312121}y_{13}+D_{13121213}y_{31})-(3\leftrightarrow 2)],$
with $X_{1}$ given by Eq. (42). The noise PSD of $T_{2}$ [71] at low frequencies is also significantly enhanced relatively to the spectrum $S_{T_{2}}=4\sin^{2}(4\pi fL)\,S_{T_{1}}$ in the SEA, where
$\frac{S_{T_{1}}}{\sin^{2}(2\pi fL)}=32\sin^{2}(\pi fL)\left[4\sin^{2}(\pi fL)\,S_{\text{acc}}+\,S_{\text{oms}}\right].$
For concreteness, we consider the soft displacement-memory signal at LISA, using the realistic orbital data provided by LISAanalysistools [57] (still adopting the static-detector approximation). Fig. 26 shows the result for $\{\phi,\theta,\psi\}=\{0.32\pi,0.52\pi,0.25\pi\}$ at $t=0.11\,\text{yr}$ and $t=0.5\,\text{yr}$ (in the time coordinate used for the orbital data), with $\{L_{23},L_{13},L_{12}\}\approx\{0.99,0.98,0.97\}L$ and $L=2.5\,\text{Gm}$. As in the A and E channels, $|\tilde{T}_{2}|/f^{2}$ also tends to a constant value at low frequencies, which is maximized at $t=0.11\,\text{yr}$ for the chosen values of $\{\phi,\theta,\psi\}$. To obtain the corresponding result in the SEA, we define an approximate detector frame by setting $\frac{1}{3}\sum_{I}\mathbf{x}_{I}=\mathbf{x}_{3}\times\mathbf{e}_{y}=\mathbf{0}$ and $\mathbf{n}_{12}\cdot\mathbf{e}_{z}=\mathbf{n}_{13}\cdot\mathbf{e}_{z}=0$, as depicted in Fig. 25. We find that the SNR of T channel is significantly larger than that in the SEA, which, however, is still much smaller than in the A and E channels. In particular, the inclusion of T channel in the parameter estimation cannot break the degeneracy of the signal at low frequencies. Compared with the displacement memory signal with the same $\{\phi,\theta,\psi\}$ parameters, the SNR of velocity memory signal in the T channel exhibits a modest increase relative to the A and E channels, since the signal is shifted to lower frequencies while $S_{T}/S_{A}$ also decreases with $f$; nevertheless, the resulting SNR remains comparably low. We leave a thorough analysis using the full set of second-generation $\{A,E,T\}$ channels or other sets of TDI combinations for future work.
Figure 25: Positions of the SC in the approximate detector frame at $t=0.11\,\text{yr}$, shown as gray points. The colored points mark the SC positions in the SEA, as defined in Fig. 8.
Figure 26: Spectrum of $|\tilde{T}_{2}|/f^{2}$ (top panel) and the logarithmic spectral density of $\rho_{A,E,T}^{2}$ (bottom panel) for $\{\phi,\theta,\psi\}=\{0.32\pi,0.52\pi,0.25\pi\}$, in the case of displacement memory. The vertical dashed line corresponds to $f=1/(2L)$.
Appendix B Soft signals from dark radiation
Similar to gravitational radiation, memory can also appear in massless scalar and vector radiation [7]. If the scalar or vector field couples to the test masses in a GW interferometer, this memory can also lead to soft signals (or their mimickers, in the case of massive fields), [^6] Note that even in the absence of non-gravitational couplings, dark radiation generally produces a gravitational displacement memory, either null or ordinary depending on whether the field is massive. as will be shown below.
Consider a Klein–Gordon field $\phi$ and a Maxwell or Proca field $A_{a}$ sourced by point charges. The worldline action of a point particle with mass $m$, scalar charge $q_{\phi}$, and vector charge $q_{A}$ is given by
$S=\int(-m+q_{\phi}\phi)\,d\tau+\int q_{A}A_{a}dX^{a},$ (78)
[…]
$\phi(u)$ $=\int_{-\infty}^{\infty}df\,\sum\tilde{\phi}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{0}(u)$ $=\int_{-\infty}^{\infty}df\sum\tilde{A}^{0}_{l,m}(f)\,Y_{lm}\,e^{-2\pi ifu},$
$A^{\text{L}}(u)$ $\equiv\mathbf{A}\cdot\hat{\mathbf{k}}=\int_{-\infty}^{\infty}df\sum\tilde{A}_{l,m}^{\text{L}}\,{}_{-1}Y_{lm}\,e^{-2\pi ifu},$
$A$ $\equiv\mathbf{A}\cdot(\mathbf{e}_{\Theta}-i\mathbf{e}_{\Phi})=\int_{-\infty}^{\infty}df\sum\tilde{A}_{l,m}\,{}_{-1}Y_{lm}\,e^{-2\pi ifu},$
(see Sec. III for the definition of $\{R,\Theta,\Phi\}$) where $u=t-v_{g}R$, with $v_{g}(f)=\sqrt{1-m^{2}/(2\pi f)^{2}}$ being the group velocity. The idealized frequency spectra of the non-vanishing modes for the 0PN scalar dipole radiation are
$\frac{\tilde{\phi}_{1,\pm 1}}{\frac{\nu\gamma aM}{R}}=\sqrt{1-\left(\frac{n_{0}}{2\varpi}\right)^{2}}\,\sqrt{\frac{\pi}{24}}\left(\pm iH_{in}^{\prime}-\frac{\sqrt{e^{2}-1}}{e}H_{in}\right),$
where $n_{0}\equiv m/\Omega$, $\varpi\equiv\pi f/\Omega$, $\gamma\equiv q_{2}/m_{2}-q_{1}/m_{1}$, and $H_{in}^{\prime}=(H_{in-1}-H_{in+1})/2$. The corresponding dipolar scalar memory for $n_{0}=0$ is given by $4\pi R\Delta\phi=\nu M\gamma\hat{\mathbf{k}}\cdot\Delta\mathbf{v}$, with $\Delta\mathbf{v}=-(2a\Omega/e)\mathbf{e}_{X}$ for the hyperbolic orbit. The 0PN dipolar vector waveform can be obtained from the scalar waveform as $\tilde{A}^{0}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}$, $\tilde{A}^{\text{L}}_{1,\pm 1}=\tilde{\phi}_{1,\pm 1}/v_{g}$ and $\tilde{A}_{1,\pm 1}=\sqrt{2}\,\tilde{\phi}_{1,\pm 1}/v_{g}$, with the replacement $q_{\phi}\to q_{A}$. The spectra are shown in Fig. 27 for different values of $n_{0}$. Even for a massive scalar field, the spectrum in the intermediate-frequency band can mimic a soft waveform for sufficiently small $n_{0}$.
Figure 27: Spectra of 0PN dipolar scalar waveforms from a charged binary on a Keplerian hyperbolic orbit.
Now we discuss the interferometer response to scalar and vector waves, assuming that the test masses are also described by the worldline action (78) (which are considered, e.g., in [98]). The acceleration of a nonrelativistic test mass induced by weak scalar and vector waves is $\mathbf{F}_{*}\approx\frac{q_{\phi}[\nabla\phi+(\dot{\mathbf{X}}\cdot\nabla\phi+\dot{\phi})\dot{\mathbf{X}}]+q_{A}[-\nabla A^{0}-\dot{\mathbf{A}}+\dot{\mathbf{X}}\times(\nabla\times\mathbf{A})]}{m-q_{\phi}\phi}\approx\frac{q_{\phi}}{m}\nabla\phi-\frac{q_{A}}{m}\left(\nabla A^{0}+\dot{\mathbf{A}}\right)$, where we neglect terms dependent on the test mass velocity $\dot{\mathbf{X}}$. For a plane wave with frequency $\omega=2\pi f$ in the direction $\hat{\mathbf{k}}$, the induced velocity of the $I$-th test mass at $\mathbf{x}_{I}$ is $\mathbf{v}_{I}(t)\equiv\int^{t}dt^{\prime}\,\mathbf{F}_{*}(t^{\prime},\mathbf{x}_{I})\approx-\frac{q_{\phi}}{m}v_{g}\phi\,\hat{\mathbf{k}}+\frac{q_{A}}{m}\left(v_{g}A^{0}\,\hat{\mathbf{k}}-\mathbf{A}\right)$, which leads to the one-way signal:
$y_{rs}(t)\approx-\mathbf{n}_{rs}\cdot[\mathbf{v}_{r}(t)-\mathbf{v}_{s}(t-L)].$ (79)
For a plane wave, the idealized spectrum $\tilde{y}_{rs}(f)$ in the static-detector approximation depends linearly on $\tilde{\phi}$ in the scalar case and on $\{\tilde{A}^{0},\tilde{A}^{\text{L}},\tilde{A}\}$ in the vector case. Soft signals can thus arise from the scalar or vector “velocity memory”, which differ from those due to nonrelativistic waves [98, 97, 95] and, in principle, are also distinguishable from the soft signals of GW memory. The same holds for potentially observable signals from the axion–photon coupling in an axion field $\phi$, for which the one-way signal is independent of $\mathbf{n}_{rs}$ [95].